### v1

# PrivaDE: Privacy-preserving Data Evaluation for Blockchain-based Data Marketplaces

Abstract.
Evaluating the relevance of data is a critical task for model builders seeking to acquire datasets that enhance model performance. Ideally, such evaluation should allow the model builder to assess the utility of candidate data without exposing proprietary details of the model. At the same time, data providers must be assured that no information about their data—beyond the computed utility score—is disclosed to the model builder.
In this paper, we present PrivaDE, a cryptographic protocol for privacy-preserving utility scoring and selection of data for machine learning. While prior works (e.g., (Qian et al., 2024; Song et al., 2021)) have proposed data evaluation protocols, our approach advances the state of the art through a practical, blockchain-centric design. Leveraging the trustless nature of blockchains, PrivaDE enforces malicious-security guarantees and ensures strong privacy protection for both models and datasets. To achieve efficiency, we integrate several techniques—including model distillation, model splitting, and cut-and-choose zero-knowledge proofs—bringing the runtime to a practical level. Furthermore, we propose a unified utility scoring function that combines empirical loss, predictive entropy, and feature-space diversity, and that can be seamlessly integrated into active-learning workflows. Evaluation shows that PrivaDE performs data evaluation effectively, achieving online runtimes within 15 minutes even for models with millions of parameters.
Our work lays the foundation for fair and automated data marketplaces in decentralized machine learning ecosystems.
conference: ; ;
1. Introduction
As machine learning (ML) systems proliferate and generate substantial value, concerns have grown over large-scale, uncompensated data collection and the need to return value to original contributors. Collaborative learning frameworks (Sim et al., 2020), in which multiple parties pool data via a decentralized ledger, offer one remedy: trainers publish data requests, contributors opt in, and compensation for contributions is transparently recorded and fairly allocated. Despite the rise of blockchain-based data and ML marketplaces (e.g., (Rao, [n. d.])), rigorous mechanisms for data evaluation remain largely overlooked.
The central challenge of data evaluation in collaborative learning is to select high-quality samples that maximally improve model performance while preserving the confidentiality of both data and model assets. Active learning (Dasgupta, 2011) addresses selection by querying the most informative examples—e.g., by uncertainty, expected loss, or diversity—thereby reducing annotation cost. Yet in a decentralized, privacy-sensitive marketplace, neither contributors nor model owners are willing to reveal raw features, labels, or model parameters prior to payment. It is therefore essential to design cryptographically secure protocols that compute data-utility scores under strong confidentiality guarantees and ensure correctness in a trustless environment.
In this paper, we study how to design a secure protocol that enables a model owner (MO) and a data owner (DO) to jointly compute a data-utility score $\phi(M,D)$ for any model $M$ and dataset $D$, with practical applicability to blockchain-based data marketplaces. This score can subsequently serve as the basis for negotiation or automated payment for data acquisition in such marketplaces.
Our Contributions. To the best of our knowledge, prior work (e.g., (Song et al., 2021; Qian et al., 2024; Tian et al., 2022)) either operates in the semi-honest model or relies on a trusted third party (TTP) for most protocol steps. PrivaDE targets deployment in blockchain-based data marketplaces under realistic assumptions: we work in the malicious-adversary setting, where either the MO or the DO may deviate arbitrarily, and we require only minimal TTP involvement limited to setup and preprocessing. Both the model $M$ and the dataset $D$ are treated as private assets and are protected throughout the protocol.
PrivaDE considers the realistic setting where the DO holds a large dataset $D$, reflecting common scenarios in which a contributor possesses many samples. Our protocol enables the MO to obtain a dataset-level score $\phi(M,D)$ while minimizing passive information leakage by requiring only a representative subset for scoring. The scoring procedure is designed so that the subset-derived score is representative of the full dataset.
We introduce several design techniques (Section 4) that bring PrivaDE, under these strict assumptions, to practical runtimes. In particular, we combine model distillation, model splitting, and cut-and-choose ZKP to balance efficiency against potential privacy leakage.
We further adapt PrivaDE to a blockchain-based data marketplace. Smart contracts deter misuse and enforce fairness in MPC by penalizing uncooperative parties via escrow slashing. The protocol can also be integrated with the data transaction to streamline the end-to-end marketplace workflow.
Finally, we prove security in the simulation-based paradigm and demonstrate practicality through extensive experiments (Section 7) on real-world datasets and standard model architectures. Results show that PrivaDE can be flexibly tuned to achieve reasonable runtimes (as low as 45 seconds online for MNIST with LeNet), and that the scoring protocol is robust, effectively identifying high-performance datasets.
An overall design of PrivaDE can be seen in figure 1.
Figure 1. Overall Design of PrivaDE. Alice first prepares the model and Bob prepares the data, then they conduct secure inference together. The inference result is then used to decide a score on the dataset.
2. Preliminaries
2.1. Machine learning
In machine learning, we compute a model $h$ from a training set $D=\{(x_{i},y_{i})\}_{i=1}^{n}$, where each $x_{i}$ has label $y_{i}$ (sometimes $z_{i}=(x_{i},y_{i})$). An algorithm $A$ searches a hypothesis space $\mathcal{H}$ for $h$ minimizing empirical loss $L_{D}(h)=\frac{1}{n}\sum_{i=1}^{n}\ell\bigl(h(x_{i}),y_{i}\bigr)$, where $\ell$ quantifies the error between prediction $y^{\prime}=h(x)$ and true $y$. Common choices include squared loss $\ell(y^{\prime},y)=(y^{\prime}-y)^{2}$ for regression and zero-one loss $\ell(y^{\prime},y)=\mathbf{1}[y^{\prime}\neq y]$ for classification. For probabilistic outputs $\vec{y}^{\prime}$, cross-entropy $\ell_{\mathrm{CE}}(\vec{y}^{\prime},\vec{y})=-\sum_{i=1}^{d}\vec{y}^{i}\ln\vec{y}^{\prime i}$ is typical, where $\vec{y}$ is a one-hot vector. We denote models by $M$ and, when parameterized by $\theta$, by $M_{\theta}$.
2.2. Cryptographic protocols
This section reviews commitment schemes, zero-knowledge proofs (ZKPs), and multi-party computation (MPC). We use $\lambda$ as the security parameter and assume familiarity with computational indistinguishability and negligible functions $\nu$.
2.2.1. Commitment schemes
Informally, a commitment scheme is a tuple $\Gamma=(\mathsf{Setup},\allowbreak\mathsf{Commit},\mathsf{Open})$ of $\mathsf{PPT}$ algorithms where:
  • $\mathsf{Setup}(1^{\lambda})\rightarrow\mathsf{pp}$ takes security parameter $\lambda$ and generates public parameters pp;
  • $\mathsf{Commit}(\mathsf{pp};m)\rightarrow(\mathtt{com},r)$ takes a secret message $m$, outputs a public commitment $Com$ and a randomness $r$ used in the commitment.
  • $\mathsf{Open}(\mathsf{pp},\mathtt{com};m,r)\rightarrow b\in\{0,1\}$ verifies the opening of the commitment $C$ to the message $m$ with the randomness $r$.
We require two main security properties: hiding and binding. The hiding requires that no $\mathsf{PPT}$ adversary, upon receiving the output of $\mathtt{com}$, can figure out what message is committed in $\mathtt{com}$. Binding requires that $\mathtt{com}$ can be opened only to a specific message. We refer to App. B.1 for the formal definition of commitment.
2.2.2. Merkle Tree Commitments
A Merkle tree commitment uses a collision-resistant hash function $H$ to commit to a vector $(m_{1},\ldots,m_{n})$ by hashing leaves and iteratively hashing pairs up a binary tree; the commitment is the root, and an opening is the $O(\log n)$-length authentication path for an index $i$.
A merkle tree commitment is a tuple of PPT algorithms $\Gamma=(\mathsf{mtSetup},\mathsf{mtCommit},\mathsf{mtOpen},\mathsf{mtVerify})$ where:
  • $\mathsf{mtSetup}(1^{\lambda})\to\mathsf{pp}$: generates public parameters $\mathsf{pp}$./
  • $\mathsf{mtCommit}(\mathsf{pp};m_{1},\ldots,m_{n})\to(\mathtt{com},\mathsf{\vphantom{p}st})$: compute leaf and internal hashes; output root $\mathtt{com}$ and state $\mathsf{\vphantom{p}st}$.
  • $\mathsf{mtOpen}(\mathsf{pp},\mathsf{\vphantom{p}st};i)\to\pi_{i}$: output the sibling-hash path authenticating position $i$.
  • $\mathsf{mtVerify}(\mathsf{pp},\mathtt{com};i,m_{i},\pi_{i})\to b$: recompute the root from $(i,m_{i},\pi_{i})$ and accept iff it equals $\mathtt{com}$.
The position-binding property (no two different $m$ open at the same $i$) holds under collision resistance of $H$. We refer to App. B.2 for the formal definition.
2.2.3. Zero-knowledge proofs
A non-interactive zero-knowledge proof-of-knowledge system (NIZKPoK) for an NP-language $L$ with the corresponding relation $\mathcal{R}_{L}$ is represented by a tuple of algorithms $\Pi=(\mathsf{Setup},\mathcal{P},\mathcal{V})$, where:
  • $\mathsf{Setup}(1^{\lambda})\rightarrow\mathtt{CRS}$ takes as the input a security parameter $\lambda$. It outputs a common reference string $\mathtt{CRS}$.
  • $\mathcal{P}(\mathtt{CRS},x,w)\rightarrow\pi$ takes as the input $\mathtt{CRS}$, the statement $x$ and the witness $w$, s.t. $(x,w)\in\mathcal{R}_{L}$. It outputs the proof $\pi$.
  • $\mathcal{V}(\mathtt{CRS},x,\pi)\rightarrow b\in\{0,1\}$ takes as the input $\mathtt{CRS}$, $x$ and $\pi$. It outputs 1 to accept and 0 to reject.
Informally, a NIZK prof is secure if it enjoys the properties of completeness, zero-knowledge and binding. Completeness simply requires that when the over and the verifier are honest the verifier always accepts the proof (i.e., it returns 1). Zero-knowledge requires that even when the verifier is malicious, the proof $\pi$ generated from a prover will leak no information about the witness $w$. Soundness requires that the verifier always rejects when receiving proofs for false statements. We refer to App. B.3 for the formal definition.
2.2.4. Multi-party computation
We prove that our results are secure under the simulation-based paradigm of secure Multi-Party Computation (MPC). In secure multiparty computation (MPC) multiple parties aim to compute a function $f$ together while keeping their inputs private. The goal is to ensure the result is correct while revealing no extra information. We say that a protocol is secure if whatever can be inferred by an adversary attacking the protocol can also be inferred by an adversary attacking a so-called ideal-world. In this, there is an ideal trusted third party that receives the inputs of the parties evaluates the function $f$, and returns the output of the computation back to the parties
In short, a protocol is secure if a real-world adversary cannot gain an advantage over what they would learn in the ideal-world setting (that trivially captures the best possible security guarantee). When defining the security of a protocol $\Pi$, we will say that $\Pi$ securely realizes the function $f$, if $\Pi$ guarantees the above. In this paper we focus on the two-party setting, considering both security against semi-honest and malicious adversaries. In the former, the corrupted parties that attack the real-world protocol are assumed to follow the protocol description (but they can try to infer information from the honest parties’ inputs). In the malicious case, instead, parties can deviate arbitrarily from what the protocol prescribes. In App. B.4 we provide a formal definition of secure MPC and its variants.
2.3. Blockchain and Smart Contracts
A blockchain is a decentralized, append-only ledger that stores an ordered sequence of blocks; each block includes transactions, a timestamp, and the previous block’s hash, yielding immutability and tamper-evidence (Nakamoto, 2008; Bonneau et al., 2015). A smart contract is a program replicated and executed by blockchain nodes that enforces state transitions and agreements without intermediaries, enabling programmable, trustless transactions (Buterin, 2014; Wood, 2014).
PrivaDE targets an EVM-compatible chain, which provides a secure, verifiable substrate for data exchange and computation. We use the chain to:
  • Commit inputs. Data contributions and model parameters are posted as cryptographic commitments, creating a binding, immutable audit trail.
  • Enforce MPC fairness. Contracts escrow deposits and impose penalties for deviations, discouraging aborts and enforcing compliance during MPC.
  • Enable atomic payment. Transfers to data contributors trigger only on verifiable completion (e.g., successful checks/proofs), preventing unilateral defection.
  • Support extensibility. The same infrastructure integrates with off-chain storage (e.g., IPFS) for availability and scalability.
3. Problem Description and Threat Model
Consider a scenario in which Alice owns a model $M$ trained on her dataset $D_{A}$ and seeks additional data to further improve it. Bob holds a relevant dataset $D_{B}$ that could enhance $M$ when combined with $D_{A}$ (i.e., $D_{A}\cup D_{B}$). Bob seeks fair compensation for his contribution, while Alice aims to pay in proportion to the improvement attributable to $D_{B}$. We assume $D_{B}$ is authentic and pertinent to Alice’s task; authenticity can be attested, for example, via signatures by the original data source or a trusted authority (Sec. 6).
Our objective is to compute a score $\phi=\mathsf{Score}(M,D_{B})$ that quantifies the improvement $D_{B}$ affords to $M$, approximating the net gain in model performance (e.g., increased accuracy on a held-out test set). In pursuit of this goal, we address two coupled questions:
  1. (1)
    What criterion should define the scoring function $\mathsf{Score}$?
  2. (2)
    How can we compute $\mathsf{Score}$ securely while protecting both Alice’s and Bob’s private data?
The challenge is to devise a criterion that reliably reflects true improvement—ordinarily requiring retraining—while remaining simple enough for efficient secure computation within practical runtimes. Consequently, we co-design the scoring rule and the secure protocol to yield a feasible and efficient scoring system.
We assume Bob contributes a larger dataset ($|D_{B}|>1$), reflecting practical scenarios such as a company representative contributing bulk records (e.g., sales data). To facilitate deployment on blockchain-based systems, PrivaDE is maliciously secure: it executes in the presence of a PPT adversary $\mathcal{A}$ who may arbitrarily deviate from the protocol, with the identity of the corrupted party fixed prior to execution. The protocol protects the honest party by detecting any deviation and safely aborting. PrivaDE requires no trusted party beyond the setup and preprocessing stages.
PrivaDE assumes the model architecture is public, while the model weights $\theta$ are Alice’s private input. Bob’s private input is the dataset $D_{B}$, formally represented by the sequences $(\overrightarrow{x},\overrightarrow{y})$. The $i$-th sample is $(x_{i},y_{i})$, the $i$-th entries of $\overrightarrow{x}$ and $\overrightarrow{y}$.
The primary goal of our protocols is to protect $\theta$ and $(\overrightarrow{x},\overrightarrow{y})$. For efficiency, we allow carefully bounded, protocol-specified leakage, supported by prior work and experiments, which does not reveal the full model $\theta$ or any individual sample $(x_{i},y_{i})$. The remaining, unleaked components are proven secure via formal, simulation-based arguments; see Sec. 5 for definitions and proofs.
4. Practical Optimizations for PrivaDE
In this section, we describe several design techniques that innovates PrivaDE and makes PrivaDE practical.
4.1. Model Distillation
Model distillation for modern neural networks, introduced by Hinton et al. (Hinton et al., 2015), is widely used to reduce model size and inference cost. The original (teacher) model guides training of a smaller (student) model, typically by matching the teacher’s logits.
Distillation is also common in secure inference (e.g., (Pereteanu et al., 2022)) to lower protocol runtime. Empirical studies (Romero et al., 2015; Cho and Hariharan, 2019) show that students with over $10\times$ fewer parameters can retain accuracy comparable to the teacher, depending on task and architecture.
In PrivaDE, we include an optional, local distillation step by the model owner prior to data evaluation. The distilled model is chosen to balance accuracy and size reduction. Beyond improving efficiency, distillation can also reduce leakage: due to limited capacity, a small student cannot faithfully encode all teacher information (Cho and Hariharan, 2019). We therefore deliberately select a student size that tolerates a minor accuracy loss in exchange for substantial runtime savings and reduced leakage. In Section 7, we show that a reduced model can still support effective data evaluation and produce results comparable to the original model.
4.2. Split Model
Figure 2. Conceptual diagram of split model in PrivaDE: Split A evaluated in a secure 2PC; split B evaluated by the data owner; split C evaluated by the model owner.
Most ML models, particularly neural networks, are compositions of layers: the output of one layer feeds the next until the final layer. This structure enables split learning (Vepakomma et al., 2018), which partitions a network at a cut layer. In a typical setup, the data owner trains the front segment and the model owner trains the back segment, coordinating forward and backward passes to reduce data exposure.
In secure inference, this idea yields split inference (e.g., (Pereteanu et al., 2022; Zhang et al., 2025a)): the front segment is executed within a secure protocol, while the remaining layers run in the clear. This reduces the cryptographic workload and thus latency, while limiting input leakage to the (carefully chosen) intermediate activations at the cut layer.
In PrivaDE, we also employ split-model inference, but propose a more sophisticated three-way split that factorizes $M$ as $C\circ B\circ A$ (Figure 2):
  1. (1)
    Block $A$. Contains the initial layers up to (and including) the first activation. We insert a secret mixer that permutes $A$’s outputs; this mixer can be an invertible linear layer or an invertible $1\times 1$ convolution that mixes channels (see (Kingma and Dhariwal, 2018)). Block $A$ is evaluated via a maliciously secure 2PC: Alice supplies weights, Bob supplies inputs. The output of $A$ is given only to Bob.
  2. (2)
    Block $B$. Comprises intermediate layers. The first layer of $B$ incorporates the inverse mixer to undo $A$’s permutation. The BBCC cut is chosen by the C2PI boundary-search procedure (Zhang et al., 2025a). Bob executes $B$ locally after receiving the corresponding weights from Alice.
  3. (3)
    Block $C$. Contains the remaining layers up to the output. Alice executes $C$ after receiving the activations produced by $B$ from Bob.
The BBCC split follows prior work aiming to curb input leakage while reducing runtime. We adopt the C2PI boundary search (Zhang et al., 2025a), which probes candidate cut layers by attempting input reconstruction from the cut-layer activations; for images, a split is accepted when the structural similarity index with the original input falls below $0.3$.
The AABB split further minimizes 2PC overhead by restricting cryptographic computation to the minimal front segment (up to the first activation), leaving the remainder in $B$. Because $B$’s weights are sent in the clear to Bob, $A$ prevents Bob from obtaining a self-sufficient front model that would yield reusable embeddings on arbitrary inputs.
Finally, we attach zero-knowledge proofs to the evaluations of $B$ and $C$ to enforce correctness and deter cheating. See Sec. 5 for protocol details.
4.2.1. Leakage of model $B$ to Bob:
Yokinski et al. (Yosinski et al., 2014) show that later layers tend to be more task-specific, whereas early layers capture generic features. Because neither the weights of $C$ nor the model outputs are revealed, Bob cannot reconstruct the full task model. However, $B\circ A$ may still yield useful embeddings and thus be valuable for fine-tuning on related domains; accordingly, we must bound the transferability of $B$.
To repurpose the model, Bob must:
  1. (1)
    Reconstruct $A$ to map raw inputs into $B$’s input space; and
  2. (2)
    Apply transfer learning to $B\circ A$ for other tasks.
For reconstructing $A$, Bob only observes his inputs and the corresponding outputs of $A$. We assume each contributor executes at most $k$ queries through $A$. A secret mixer (which can be different on different runs) applies a random invertible permutation to $A$’s outputs, which thwarts simple plug-and-play (attaching public pretrained fronts to $B$) and impedes collusion by multiple contributors pooling I/O pairs for extraction.
Empirical reports (Orekondy et al., 2019) suggest successful extraction often requires on the order of $10^{4}$ queries; thus, security improves when $k\ll 10^{4}$. To quantify robustness at budget $k$, we propose a reconstruction-under-budget experiment: randomly sample $\{x_{i}\}_{i=1}^{k}$ from $D_{A}$ and train $A^{\prime}$ (a re-initialized copy of $A$) by minimizing
$\min_{\theta}\;\frac{1}{k}\sum_{i=1}^{k}\big\|A^{\prime}(x_{i};\theta)-u_{i}\big\|_{2}^{2}\;+\;\lambda\,\big\|B\big(A^{\prime}(x_{i};\theta)\big)-B(u_{i})\big\|_{2}^{2},$
with $\lambda>0$ weighting feature consistency through $B$. After training, fix $B$ and $C$, replace $A$ by $A^{\prime}$, and report top-1 accuracy on the standard test set. The primary metric is
$\Delta\mathrm{Acc}\;=\;\mathrm{Acc}\big(C\!\circ\!B\!\circ\!A^{\prime}\big)\;-\;\mathrm{Acc}\big(C\!\circ\!B\!\circ\!A\big).$
A split is deemed robust if $\Delta\mathrm{Acc}$ is substantially negative at practical $k$, indicating that $A$ is not reconstructable and $B$ is not easily reusable under the stated constraints (cf. the privacy criterion in Zhang et al., 2025a).
We also assess the transferability of $B\circ A$ (hypothetically assuming $A$ is learned). Following You et al. (You et al., 2021), we compute LogME, the log marginal evidence that a Bayesian linear regressor explains labels from features. Concretely, we pass labeled images through $B(A(x))$ to obtain features and take the resulting LogME value as a transferability score; lower scores indicate poorer linear transfer and thus reduced leakage risk.
Alice decides by herself if the leakage is acceptable based on $\Delta\mathrm{Acc}$ and the LogME score by comparing it to baselines of other models in the marketplace.
4.2.2. Leakage of activations $B(A(x))$ to Alice:
We quantify input leakage by reconstruction quality, training an inverse model that maps $B(A(x))$ back to $x$. For images, we adopt DINA (Zhang et al., 2025a) and report the average structural similarity index (SSIM) over a test set; we accept the split when the mean SSIM <0.3<0.3, indicating low-fidelity reconstructions.
4.3. Representative Set Selection
Because Bob contributes a dataset $D_{B}$, scoring every point in $D_{B}$ is costly and increases leakage: repeated inferences can expose information about Alice’s model and Bob’s data distribution.
To mitigate this, our protocol requires Bob to extract a much smaller representative subset $D_{R}\subset D_{B}$, and then jointly and securely evaluate Alice’s model only on $D_{R}$. To ensure correctness even if Bob is malicious, we enforce honest sampling via a cut-and-choose subroutine (Sec. 4.3.2). Similar representative-sampling strategies (without security) are standard in batch active learning (Sener and Savarese, 2018) and related ML tasks (Mirzasoleiman et al., 2013).
We formalize representativeness as follows:
Definition 4.1 (Representativeness of $D_{R}$ to $D_{B}$).
Let $D_{R}\subset D_{B}$, and fix a distance threshold $d>0$ and outlier tolerance $\delta\in[0,1]$. Then $D_{R}$ is $(d,\delta)$-representative if $\Pr_{x\sim D_{B}}\!\bigl[\exists\,x^{\prime}\in D_{R}:\|x-x^{\prime}\|<d\bigr]\geq 1-\delta$, where the probability is with respect to the empirical distribution over $D_{B}$.
We write $\mathcal{R}(D_{B})$ for any algorithm that, given $|D_{B}|=n>k$, returns $D_{R}\subset D_{B}$ of size $k$. Common methods include $k$-center clustering and submodular maximization (Nemhauser et al., 1978; Mirzasoleiman et al., 2013). The parameters $(k,d,\delta)$ are chosen jointly by Alice and Bob per application.
4.3.1. Dimension Reduction.
The selection step involves pairwise distances as in Def. 4.1, which becomes expensive in $\mathbb{R}^{d}$ as $d$ grows. We therefore apply random linear projections (Bingham and Mannila, 2001):
$x^{\prime}=R\,x,\qquad R\in\mathbb{R}^{m\times d},\;m\ll d,$
which preserve distances up to $\varepsilon$-distortion by the Johnson–Lindenstrauss lemma (Johnson and Lindenstrauss, 1984). In a secure setting, Alice and Bob jointly sample $R$ via $md$ secure coin flips (Blum, 1981), deriving a shared pseudorandom seed. Bob projects his data locally and runs the representative-selection protocol on $\{x^{\prime}\}$, reducing computation and communication while retaining scoring fidelity.
4.3.2. Challenge Protocol for Representative Set Selection
A naive way to securely implement the representative-set algorithm $\mathcal{R}$ is via a maliciously secure two-party computation (2PC). However, because state-of-the-art clustering often relies on costly Euclidean distances, this approach incurs substantial overhead.
As an alternative, we introduce the challenge protocol $\mathsf{CP}$. Bob commits each point $x_{i}\in D_{B}^{x}$ ($D_{B}^{x}:=\{x\mid(x,y)\in D_{B}\}$) via commitments $\mathtt{com}_{1},\dots,\mathtt{com}_{n}$, and publishes the representative set $D_{R}$ by revealing its indices $I_{R}:=\{i|x_{i}\in D_{R}^{x}\}$.
Alice selects a random index $i\in[n]$ and challenges Bob to prove, in zero knowledge, that the committed point $x_{i}$ is well represented by $D_{R}$. Bob recovers $x_{i}$, computes $d_{i}^{\prime}=\min_{x^{r}_{j}\in D_{R}^{x}}\|x^{r}_{j}-x_{i}\|$, and:
  • if $d_{i}^{\prime}<d$, returns a ZK proof $\pi_{i}$ attesting that $\exists j,r_{i}:\ \|x^{r}_{j}-x_{i}\|\leq d\ \wedge\ \mathsf{Open}(\mathsf{pp},\mathtt{com}_{i},x_{i},r_{i})=1$;
  • otherwise replies $\mathsf{fail}_{i}$.
By repeating this challenge, Alice gains confidence that $D_{R}$ is representative of $D_{B}$. The full protocol appears in Fig. 3, and its security is established in the next section.
Parties: Model owner $P_{1}$, data owner $P_{2}$ Public parameters: Indices $I_{R}=\{i_{1},\dots,i_{k}\}$; commitments of $D_{B}^{x}$: $\mathsf{COM}_{B}=\{\mathtt{com}_{1},\ldots,\mathtt{com}_{n}\}$; commitment parameter $\mathsf{pp}$; maximum distance $d$; outlier ratio $\delta$ (1) $P_{1}$ samples a set of indices $I\subset[n]$ uniformly at random and sends it to $P_{2}$. (2) For each $j\in I$, $P_{2}$ recovers $x_{i}\in D_{B}^{x}$ and computes $d_{i}^{\prime}=\min_{x^{r}_{j}\in D_{R}^{x}}\|x^{r}_{j}-x_{i}\|$. (3) For each $i\in I$: If $d_{i}^{\prime}<d$, $P_{2}$ sends a ZK proof $\pi_{i}$ for the language $L_{\mathsf{CP}}$ $=\Bigl\{\,(\mathsf{pp},d,i,\mathsf{COM}_{B},I_{R})\ \Bigm|\ \exists\,j,x_{i},r_{i},x_{j},r_{j}\ \text{s.t.}$ $\hskip 18.49988ptj\in I_{R},\qquad\|x^{r}_{j}-x_{i}\|\leq d,$ $\hskip 18.49988pt\mathsf{Open}(\mathsf{pp},\mathtt{com}_{i},x_{i},r_{i})=1,\qquad\mathsf{Open}(\mathsf{pp},\mathtt{com}_{j},x_{j},r_{j})=1\Bigr\}.$ Otherwise, $P_{2}$ replies $\mathsf{fail}_{i}$. (4) $P_{1}$ verifies each $\pi_{i}$. If at least $(1-\delta)\,|I|$ proofs succeed, the protocol accepts; otherwise, $P_{1}$ aborts.
Figure 3. Challenge Protocol $\mathsf{CP}$ for $D_{R}$ representativeness verification.
If Alice requires $D_{R}$ to be $(d,\delta)$-representative of $D_{B}$, then running $|I|=\Theta(\ln n/\delta)$ challenges suffices to ensure, with probability at least $1-n^{-c}$, that at most a $\delta$ fraction of points are outliers after $c\ln n/\delta$ successful proofs. In the full PrivaDE protocol, $d$ and $\delta$ are pre-agreed, and $I\subset[n]$ denotes the indices Alice selects for $\mathsf{CP}$.
4.4. Audit-Style Cut-and-Choose for ML Inference
While significant research has focused on zero-knowledge proofs (ZKPs) for verifying ML inference (e.g., (Huang et al., 2022; Kang et al., 2022; Liu et al., 2021)), constructing end-to-end proofs remains memory- and computation-intensive.
Instead of requiring a full proof, we propose a lightweight protocol that detects incorrect batched inference with tunable detection probability (e.g., 80–95%) by sampling and proving only a small, random subset of local layer transitions. The protocol $\mathsf{CnCZK}$ serves as a drop-in replacement for a NIZKPoK of the following language $L_{\textsf{ML-Inf}}$. We consider two scenarios: (1) the model weights and outputs are hidden as part of the witness and the data point is public (blue); (2) the data point is hidden as part of the witness and the model weights and outputs are public (red). For each scenario, read the black text together with the text in the corresponding colour.
$L_{\textsf{ML-Inf}}=\left\{\,\begin{array}[]{l}(\mathsf{pp},{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\overrightarrow{x},\mathtt{com}_{\theta},\overrightarrow{\mathtt{com}_{y^{\prime}}}},\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\theta,\overrightarrow{\mathtt{com}_{x}},\overrightarrow{y^{\prime}}})\end{array}:\begin{array}[]{l}\exists\,({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\theta,r_{\theta},\overrightarrow{y^{\prime}},\overrightarrow{r_{y^{\prime}}}},{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\overrightarrow{x},\overrightarrow{r_{x}}})\ \text{s.t.}\\ M_{\theta}(x)=y^{\prime},\\ {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathsf{Open}(\mathsf{pp},\mathtt{com}_{\theta},\theta,r_{\theta})=1},\\ {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\forall y^{\prime}\,\mathsf{Open}(\mathsf{pp},\mathtt{com}_{y^{\prime}},y^{\prime},r_{y^{\prime}})=1},\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\forall x\,\mathsf{Open}(\mathsf{pp},\mathtt{com}_{x},x,r_{x})=1}\end{array}\right\}.$
Notation.
Let $M_{\theta}$ be a fixed $L$-layer network with parameters $\theta$ (with $\theta_{\ell}$ the weights at layer $\ell$), inputs $\{x_{i}\}_{i=1}^{N}$, intermediate activations $a_{\ell,i}$ for $\ell\in\{0,\dots,L\}$ (with $a_{0,i}=x_{i}$), and outputs $y^{\prime}_{i}=a_{L,i}$. Denote the $\ell$-th layer by $m_{\ell}(\theta_{\ell};\cdot)$, so $a_{\ell,i}=m_{\ell}(\theta_{\ell};a_{\ell-1,i})$.
The protocol $\mathsf{CnCZK}$ proceeds as follows.
Parties: Prover $P_{1}$, Verifier $P_{2}$ Public parameters: Public parameter $\mathsf{pp}$; audit sizes $m$ (checked points) and $s$ (checked layers per point); the inputs $\{x_{i}\}_{i=1}^{N}$; the model weights $\{\theta_{j}\}_{j=1}^{L}$. (1a) Model binding. The prover fixes the model parameters $\{\theta_{j}\}_{j=1}^{L}$ for the audit window and publishes a model commitment as a Merkle root $R_{\theta}$: $R_{\theta},\mathsf{\vphantom{p}st}_{m}\leftarrow\mathsf{mtCommit}(\mathsf{pp};\theta_{1},\ldots,\theta_{L})$. (1b) Data binding. The prover fixes the inputs $\{x_{i}\}_{i=1}^{N}$ for the audit window and publishes a data commitment as a Merkle root $R_{X}$: $R_{X},\mathsf{\vphantom{p}st}_{d}\leftarrow\mathsf{mtCommit}(\mathsf{pp};x_{1},\ldots,x_{N})$. (2) Compute and commit the full trace. The prover runs a full forward pass on all NN inputs, obtaining $\{a_{\ell,i}\}$ and $\{y^{\prime}_{i}\}$, and publishes Merkle commitments $\{R_{\ell}\}_{\ell}$: $R_{\ell},\mathsf{\vphantom{p}st}_{\ell}\;\leftarrow\;\mathsf{mtCommit}\big(\mathsf{pp};a_{\ell,1},\ldots,a_{\ell,N}\big)\quad\text{for all }\ell\in\{0,\ldots,L\},$ where $a_{0,i}=x_{i}$ and $a_{L,i}=y^{\prime}_{i}$. (3) Challenge sampling. The verifier samples a set $S\subseteq[N]$ of $m$ distinct datapoints (uniform without replacement). For each $x_{i}\in S$, the verifier independently samples a set $T_{i}\subseteq\{1,\ldots,L\}$ of $s$ distinct layers (uniform without replacement). (4) Prove local correctness on sampled layers. For each $x_{i}\in S$ and each $\ell\in T_{i}$, the prover produces a ZKP for the NP language $L_{\textsf{CnCZK}}$ and sends to the verifier: $L_{\textsf{CnCZK}}=\left\{\begin{array}[]{l}(\mathsf{pp},i,\ell,{\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}R_{\theta}},\\ {\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\theta_{\ell}},R_{\ell-1},R_{\ell})\end{array}:\begin{array}[]{l}\exists\,({\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\theta_{\ell},\pi_{\theta}},a_{\ell-1,i},a_{\ell,i},\pi_{\ell-1},\pi_{\ell})\ \text{s.t.}\\[2.0pt] m_{\ell}(\theta_{\ell};a_{\ell-1,i})=a_{\ell,i},\\[2.0pt] {\color[rgb]{0,0,1}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,1}\mathsf{mtVerify}(\mathsf{pp},R_{\theta},\ell,\theta_{\ell},\pi_{\ell})=1},\\[2.0pt] \mathsf{mtVerify}(\mathsf{pp},R_{\ell-1},i,a_{\ell-1,i},\pi_{\ell-1})=1,\\[2.0pt] \mathsf{mtVerify}(\mathsf{pp},R_{\ell},i,a_{\ell,i},\pi_{\ell})=1\end{array}\right\}.$ Additionally, the prover sends to the verifier a Merkle opening for $R_{0}$ to show that $x_{i}$ is included in $R_{0}$: $\pi\leftarrow\mathsf{mtOpen}(\mathsf{pp},\mathsf{\vphantom{p}st}_{0},i)$ Additionally, the prover sends to the verifier a ZKP for the input-consistency language $L_{\textsf{in}}$ to ensure the correct inputs are used: $L_{\textsf{in}}=\left\{(\mathsf{pp},R_{0},R_{X},i):\begin{array}[]{l}\exists\,x_{i},\pi_{1},\pi_{2}\ \text{s.t.}\\[2.0pt] \mathsf{mtVerify}(\mathsf{pp},R_{0},i,x_{i},\pi_{1})=1,\\[2.0pt] \mathsf{mtVerify}(\mathsf{pp},R_{X},i,x_{i},\pi_{2})=1\end{array}\right\}.$ (5) Verification and decision. The verifier checks all proofs and merkle openings. Accept if and only if all sampled checks pass; otherwise, reject.
Figure 4. Cut-and-choose Protocol $\mathsf{CnCZK}$ for model inference verification.
4.4.1. Cheating-detection probability
Because the scoring protocol aggregates all evaluated points, altering a single point has limited impact on the final score. We therefore consider an adversary that corrupts a fraction $\rho\in(0,1)$ of the $N$ inputs.
Theorem 4.2.
Consider $\mathsf{CnCZK}$ on a model with $L$ layers and $N$ inference points, auditing $m$ points and $s$ layers per audited point (both sampled uniformly without replacement). If an adversary corrupts $N\rho$ points, the probability of detecting cheating is bounded below by
$\Pr[\textsf{detect}]\;\geq\;\sum_{k=0}^{\min(N\rho,\,m)}\frac{\binom{N\rho}{k}\binom{N(1-\rho)}{\,m-k\,}}{\binom{N}{m}}\Bigl[1-\Bigl(\tfrac{L-s}{L}\Bigr)^{k}\Bigr].$
The proof appears in Appendix A.1. As an illustration, with $N=100$, $L=10$, $\rho=0.10$, $m=25$, and $s=6$, the detection probability exceeds $80\%$. This setting audits only $ms/(NL)=150/1000=15\%$ of all layer evaluations, yielding about an $85\%$ reduction relative to checking every layer of every point.
5. PrivaDE: The Data Evaluation Protocol
In this section, we describe the end-to-end workflow of PrivaDE.
5.1. Dataset Scoring Algorithm
Given a representative set $D_{R}$, the parties compute a score using three criteria—diversity, uncertainty, and loss—where diversity and uncertainty follow standard active-learning practice.
Diversity. The subset should be diverse, meaning elements of $\mathcal{S}$ are sufficiently different on average to avoid redundancy. We denote diversity by $\mathcal{D}$, instantiated, e.g., as a statistic (mean, median) of pairwise distances. Diverse-set constructions are common in coreset and batch active learning (Sener and Savarese, 2018).
Uncertainty. Uncertainty indicates potential informativeness: high uncertainty suggests the model benefits from additional training on such samples. A typical measure is the entropy of $M$’s predictive distribution; margin-based variants are also used. We write the uncertainty functional as $\mathcal{U}$.
Loss. Relevance can be captured by the loss incurred by $M$ on Bob’s data (Yoo and Kweon, 2019). If $y^{\prime}=M_{\theta}(x)$, then $\ell(y^{\prime},y)$ quantifies the discrepancy between prediction and label; assuming valid data, larger loss indicates greater potential value for training.
Let
$l$ $=\frac{1}{|D_{R}|}\;\;\sum_{\mathclap{(x,y)\in D_{R}}}\ell\!\left(M_{\theta}(x),y\right),$
$u$ $=\mathcal{U}\!\left(M_{\theta}[D_{R}]\right),$
$d$ $=\mathcal{D}\!\left(D_{B}^{x},D_{R}^{x}\right).$
The overall score is $\phi=f(l,u,d)$ for a pre-agreed aggregator $f$. Algorithm 1 summarizes the cleartext procedure, where $D_{B}^{x}=\{x\mid(x,y)\in D_{B}\}$ and $D_{B}^{y}=\{y\mid(x,y)\in D_{B}\}$.
Algorithm 1 Cleartext dataset scoring $\mathsf{Score}_{\mathsf{multi}}$
1:model $M_{\theta}$, dataset $D_{B}$, loss $\ell$, uncertainty $\mathcal{U}$, diversity $\mathcal{D}$, scoring $f$, representative sampler $\mathcal{R}$
2:score $\phi$ of $D_{B}$
[…]
4:For each $(x_{i},y_{i})\in D_{R}$, compute $y^{\prime}_{i}\leftarrow M_{\theta}(x_{i})$
[…]
7:return $\phi$
Example.
For multi-class image classification, choose cross-entropy loss $\ell_{\mathrm{CE}}$, margin-based uncertainty $u=\frac{1}{k}\sum_{i}(\max p_{i}-\text{second-max }p_{i})$ (Settles, 2009), diversity $d=$ mean pairwise $\|x-x^{\prime}\|$ over $D_{R}^{x}$, and $f(l,u,d)=\alpha l+\beta u+\gamma d$ for weights $\alpha,\beta,\gamma$.
5.2. The PrivaDE Protocol
We now give a secure realization of Algorithm 1. At a high level, the protocol proceeds in four stages:
  1. (i)
    Setup. Public parameters, keys, and any preprocessing required by the underlying secure primitives are initialized.
  2. (ii)
    Representative selection & audit. Bob computes a representative set $D_{R}$, and Alice audits it using the challenge protocol $\mathsf{CP}$ (Sec. 4.3.2) to ensure $(d,\delta)$-representativeness on sufficiently many points.
  3. (iii)
    Verified inference. The model is evaluated on $D_{R}$, and the cut-and-choose protocol $\mathsf{CnCZK}$ is used to verify sampled layer computations.
  4. (iv)
    Secure scoring. Alice and Bob invoke a maliciously secure 2PC that realizes $\mathcal{F}_{\mathsf{SubScore}}$ (Fig. 7) to compute the final score from loss, uncertainty, and diversity statistics.
Formally, we prove that the protocol realizes $\mathcal{F}_{\mathsf{Score}}$, which securely computes Algorithm 1. In particular, Alice’s input (the model weights) and Bob’s input are protected except for some intended leakage. This leakage profile is fully specified by $\mathcal{F}_{\mathsf{Score}}$ in Fig. 5.
Parameters: model $M=C\circ B\circ A$, loss $\ell$, uncertainty function $\mathcal{U}$, diversity function $\mathcal{D}$, scoring function $f$, representative set selection $\mathcal{R}$, distance threshold $d$, outlier ratio $\delta$, commitment parameter $\mathsf{pp}$ (1) $\mathcal{F}_{\mathsf{Score}}$ receives $I_{R}$ (indices of representative set) from $P_{2}$. (2) $\mathcal{F}_{\mathsf{Score}}$ receives a list of challenge indices II from $P_{1}$, and sends $I$ to $P_{2}$. (3) $\mathcal{F}_{\mathsf{Score}}$ receives $D^{x}_{B|I}$ (The set $D_{B}^{x}$ indiced by II) and $D^{x}_{R|I}$ (respective points in $D^{x}_{R}$ closest to each $D^{x}_{B|I}$) from $P_{2}$. For each $i\in I$ $(x_{i}\in D^{x}_{B|I})$, $\mathcal{F}_{\mathsf{Score}}$ checks that there exists some $x_{j}\in D_{R|I}^{x}$ such that $\|x_{i}-x_{j}\|<d$. If this check fails for at least $\delta\cdot|I|$ indices, the functionality sends $\mathsf{abort}$ to both $P_{1}$ and $P_{2}$. (4) $\mathcal{F}_{\mathsf{Score}}$ receives $\theta_{A}$ from $P_{1}$ and $D^{x}_{R}$ from $P_{2}$, then computes $\overrightarrow{a_{A}}\leftarrow A_{\theta_{A}}(D^{x}_{R})$ and sends $\overrightarrow{a_{A}}$ to $P_{2}$. (5) $\mathcal{F}_{\mathsf{Score}}$ receives $\theta_{B}$ from $P_{1}$, sends $\theta_{B}$ to $P_{2}$, then computes $\overrightarrow{a_{B}}\leftarrow B_{\theta_{B}}(\overrightarrow{a_{A}})$ and sends $\overrightarrow{a_{B}}$ to $P_{1}$ and $P_{2}$. (6) $\mathcal{F}_{\mathsf{Score}}$ receives $\theta_{C}$ from $P_{1}$, computes $\overrightarrow{y^{\prime}}\leftarrow C_{\theta_{C}}(\overrightarrow{a_{B}})$. (7) $\mathcal{F}_{\mathsf{Score}}$ receives $D^{x}_{B},D^{y}_{R}$ from $P_{2}$ and calculates the score: $l=\frac{1}{k}\sum_{i=1}^{k}\ell(y^{\prime}_{i},y_{i}),\quad u=\mathcal{U}(\{y^{\prime}_{1},\ldots,y^{\prime}_{k}\}),\quad d=\mathcal{D}(D_{B}^{x},D_{R}^{x})$, and finally $\phi\leftarrow f(l,u,d)$. It then sends $\phi$ to both $P_{1}$ and $P_{2}$.
Figure 5. Functionality $\mathcal{F}_{\mathsf{Score}}$ for computing the score of a dataset.
We are now ready to formally describe the protocol PrivaDE that realizes $\mathcal{F}_{\mathsf{Score}}$ (Fig. 5). PrivaDE uses the following tools:
  • $(\mathsf{SetupCom},\mathsf{Commit},\mathsf{Open})$ is a commitment scheme satisfying Definition B.1.
  • $(\mathsf{SetupZKP},\mathcal{P},\mathcal{V})$ is an NIZKPoK system satisfying Definition 2.2.3, defined for the NP language $L_{\textsf{ML-Inf}}$ (Sec. 4.4). We specify the protocol with end-to-end ZKPs, but in practice this can be replaced by the weaker $\mathsf{CnCZK}$ audit with tunable detection probability. Our proof-of-concept implementation uses EZKL (zkonduit, 2023).
  • $\Pi_{\mathsf{Inference}}$ is a 2-party protocol satisfying Definition 2.2.4, which securely realizes $\mathcal{F}_{\mathsf{Inference}}$ (Fig. 6) with malicious security. We instantiate this with SPDZ2k (Cramer et al., 2018) in MP-SPDZ (Keller, 2020).
  • $\Pi_{\mathsf{SubScore}}$ is a 2-party protocol satisfying Definition 2.2.4, which securely realizes $\mathcal{F}_{\mathsf{SubScore}}$ (Fig. 7); also instantiated with SPDZ2k.
Parameters. Description of the model AA; model owner $P_{1}$; data owner $P_{2}$; commitments $\mathtt{com}_{A},\overrightarrow{\mathtt{com}_{x}}$; commitment parameter $\mathsf{pp}$. (1) $\mathcal{F}_{\mathsf{Inference}}$ receives private input $(\theta_{A},r_{A})$ from $P_{1}$ and $(\overrightarrow{x},\overrightarrow{r_{x}})$ from $P_{2}$. (2) it verifies the commitments $b_{A}\leftarrow\mathsf{Open}(\mathsf{pp},\mathtt{com}_{A},\theta_{A},r_{A})\quad\text{and}\quad b_{x_{i}}\leftarrow\mathsf{Open}(\mathsf{pp},\mathtt{com}_{x_{i}},x_{i},r_{x_{i}})$ for all ii. If $b_{A}\neq 1$ or $b_{x_{i}}\neq 1$ for any $i$, it sends $\mathsf{abort}$ to both $P_{1}$ and $P_{2}$. (3) It computes $\overrightarrow{y^{\prime}}\leftarrow A(\overrightarrow{x})$, commits $\mathtt{com}_{a_{i}},r_{a_{i}}\leftarrow\mathsf{Commit}(\mathsf{pp},y_{i})$ for each $y_{i}\in\overrightarrow{y^{\prime}}$ (4) It delivers $\overrightarrow{y^{\prime}},\{\mathtt{com}_{a_{i}},r_{a_{i}}\}_{i}$, to $P_{2}$, and only $\{\mathtt{com}_{a_{i}}\}_{i}$ to $P_{1}$.
Figure 6. Functionality $\mathcal{F}_{\mathsf{Inference}}$ for maliciously secure 2PC.
Parameters. Loss $\ell$; uncertainty $\mathcal{U}$; diversity $\mathcal{D}$; aggregator $f$; model owner $P_{1}$; data owner $P_{2}$; commitment parameter $\mathsf{pp}$; public representative features $D_{R}^{x}$. (1) $\mathcal{F}_{\mathsf{SubScore}}$ receives private input $(y^{\prime}_{1},\ldots,y^{\prime}_{k},r_{y^{\prime}_{1}},\ldots,r_{y^{\prime}_{k}},\mathtt{com}_{y_{1}},\ldots,\mathtt{com}_{y_{k}})$ from $P_{1}$ and $(D_{B}^{x},y_{1},\ldots,y_{k},r_{y_{1}},\ldots,r_{y_{k}},\mathtt{com}_{y^{\prime}_{1}},\ldots,\mathtt{com}_{y^{\prime}_{k}})$ from $P_{2}$. (2) It verifies the commitments $b_{y_{i}}\leftarrow\mathsf{Open}(\mathsf{pp},\mathtt{com}_{y_{i}},y_{i},r_{y_{i}})\quad\text{and}\quad b_{y^{\prime}_{i}}\leftarrow\mathsf{Open}(\mathsf{pp},\mathtt{com}_{y^{\prime}_{i}},y^{\prime}_{i},r_{y^{\prime}_{i}})$ for all ii. If $b_{y_{i}}\neq 1$ or $b_{y^{\prime}_{i}}\neq 1$ for any $i$, it sends $\mathsf{abort}$ to both $P_{1}$ and $P_{2}$. (3) It computes the average loss $l=\frac{1}{k}\sum_{i=1}^{k}\ell(y^{\prime}_{i},y_{i}),$ the uncertainty score $u=\mathcal{U}(\{y^{\prime}_{1},\ldots,y^{\prime}_{k}\}),$ the diversity score $d=\mathcal{D}(D_{B}^{x},D_{R}^{x})$ and finally $\phi\leftarrow f(l,u,d)$. It then outputs $\phi$ to both $P_{1}$ and $P_{2}$.
Figure 7. Functionality $\mathcal{F}_{\mathsf{SubScore}}$ for maliciously secure 2PC.
The Setup phase. The setup phase is an offline stage before the protocol between Alice and Bob begins. In this phase:
  • Alice and Bob communicate with a trusted third party to generate setup parameters $\mathsf{pp}\leftarrow\mathsf{SetupCom}(1^{\lambda})$ and $\mathtt{CRS}\leftarrow\mathsf{SetupZKP}(1^{\lambda})$.
  • Bob verifies the authenticity of his dataset $D_{B}$ with a data authority, which then issues a signature attesting to the authenticity of $D_{B}$.
  • Alice optionally performs model distillation locally (no verification required), then interacts with a model authority to split the model $M$ into $C\circ B\circ A$, with the authority validating that the DINA reconstruction SSIM is <0.3<0.3 (Sec. 4.2.2).
This is the only phase that requires a trusted third party. We then proceed to the main protocol, described in Fig. 8.
Parties: Model owner $P_{1}$, data owner $P_{2}$ Private parameters: Dataset $D_{B}$ (held by $P_{2}$); model weights $\theta_{A},\theta_{C}$ (held by $P_{1}$) Public parameters: Commitment parameter $\mathsf{pp}$, representative set size $k$, distance threshold $d$, outlier ratio $\delta$, model architecture $M$, number of challenges $m$, model weights $\theta_{B}$ Stage 0: Commitments (1) $P_{1}$ computes $(\mathtt{com}_{A},r_{A})\leftarrow\mathsf{Commit}(\mathsf{pp},\theta_{A})$ and $(\mathtt{com}_{C},r_{C})\leftarrow\mathsf{Commit}(\mathsf{pp},\theta_{C})$, and sends $\mathtt{com}_{A},\mathtt{com}_{C}$ to $P_{2}$. (2) For each $(x_{i},y_{i})\in D_{B}$, $P_{2}$ computes $(\mathtt{com}_{x_{i}},r_{x_{i}})\leftarrow\mathsf{Commit}(\mathsf{pp},x_{i})$ and $(\mathtt{com}_{y_{i}},r_{y_{i}})\leftarrow\mathsf{Commit}(\mathsf{pp},y_{i})$, and sends $\mathtt{com}_{x_{i}},\mathtt{com}_{y_{i}}$ to $P_{1}$. Stage 1: Representative Set Selection (3) $P_{2}$ runs the representative set selection algorithm $\mathcal{R}$ locally to select a subset $D_{R}$ of size $k$ from $D_{B}$. It then computes $d^{\prime}$ as the $(1-\delta)$-th percentile of the pairwise distances $\Big\{\min_{b\in D_{R}^{x}}\|a-b\|:a\in D_{B}^{x}\setminus D_{R}^{x}\Big\}$. If $d^{\prime}>d$, $P_{2}$ aborts the protocol. Otherwise, he sends indices $I_{R}$ to $P_{1}$. (4) $P_{1}$ and $P_{2}$ runs the challenge protocol $\mathsf{CP}$, where $P_{1}$ supplies a list of challenge indices II of size m. If the challenge protocol fails, $P_{1}$ aborts. Stage 2: Model Inference (5) $P_{1}$ and $P_{2}$ run the 2PC protocol $\Pi_{\mathsf{Inference}}$ to evaluate block AA on $D_{R}^{x}$. $P_{1}$ inputs $(\theta_{A},r_{A})$; $P_{2}$ inputs $D_{R}^{x}$ (and the corresponding opening randomness). $P_{2}$ receives the activations $\overrightarrow{a_{A}}=\{a_{A,i}\}_{i}$ and per-sample commitments $\{(\mathtt{com}_{a_{i}},r_{a_{i}})\}_{i}$; both parties obtain $\{\mathtt{com}_{a_{i}}\}_{i}$. (6) $P_{2}$ locally computes $\overrightarrow{a_{B}}\leftarrow B_{\theta_{B}}(\overrightarrow{a_{A}})$. He generates a ZK proof $\pi_{B}$ for the language $L_{\textsf{ML-Inf}}$ (public weights, hidden inputs) with witness $(\overrightarrow{a_{A}},\{r_{a_{i}}\}_{i})$, and sends $(\overrightarrow{a_{B}},\pi_{B})$ to $P_{1}$. (7) $P_{1}$ verifies $\pi_{B}$; if verification fails, she aborts. (8) $P_{1}$ computes $\overrightarrow{y^{\prime}}\leftarrow C_{\theta_{C}}(\overrightarrow{a_{B}})$ and for each $i$ computes $(\mathtt{com}_{y^{\prime}_{i}},r_{y^{\prime}_{i}})\leftarrow\mathsf{Commit}(\mathsf{pp},y^{\prime}_{i})$. She then produces a ZK proof $\pi_{C}$ for $L_{\textsf{ML-Inf}}$ (hidden weights, public inputs) with witness $(\theta_{C},r_{C},y^{\prime}_{1},\ldots,y^{\prime}_{k},r_{y^{\prime}_{1}},\ldots,r_{y^{\prime}_{k}})$, and sends $\{\mathtt{com}_{y^{\prime}_{i}}\}_{i=1}^{k}$ and $\pi_{C}$ to $P_{2}$. (9) $P_{2}$ verifies $\pi_{C}$; if verification fails, he aborts. Stage 3: Dataset Scoring (10) $P_{1}$ and $P_{2}$ run the maliciously secure 2PC protocol $\Pi_{\mathsf{SubScore}}$. $P_{1}$ inputs $\big(y^{\prime}_{1},\ldots,y^{\prime}_{k},\,r_{y^{\prime}_{1}},\ldots,r_{y^{\prime}_{k}},\,\mathtt{com}_{y^{\prime}_{1}},\ldots,\mathtt{com}_{y^{\prime}_{k}}\big)$; $P_{2}$ inputs $\big(D_{B}^{x},\,y_{1},\ldots,y_{k},\,r_{y_{1}},\ldots,r_{y_{k}},\,\mathtt{com}_{y_{1}},\ldots,\mathtt{com}_{y_{k}}\big)$. Both parties receive the score $\phi$..
Figure 8. PrivaDE for secure dataset scoring.
Theorem 5.1.
PrivaDE securely realizes $\mathcal{F}_{\mathsf{Score}}$ with malicious security.
The proof of Theorem 5.1 will be provided in Appendix C.1.
6. Blockchain Integration for PrivaDE
In this section, we describe how PrivaDE can be integrated into a blockchain-based decentralized data marketplace. This is a proof-of-concept; detailed implementation aspects are out of scope.
Using a blockchain brings four major advantages. First, commitments and data requests can be posted on-chain, enhancing transparency and guaranteeing input immutability even before evaluation begins. Second, the fairness of MPC within PrivaDE—not typically guaranteed—can be enforced by slashing escrow from dishonest parties. Third, ZKP generation in $\mathsf{CP}$ and $\mathsf{CnCZK}$ can be tied to on-chain payments, deterring overuse and improving efficiency. Finally, data transactions can be integrated with evaluation so that, for example, when the score produced by PrivaDE exceeds a threshold, the transaction is triggered atomically.
Figure 9. Decentralized Data Marketplace Workflow with PrivaDE integration. Commitments and protocol results are posted on-chain for transparency, and data transaction is seamlessly integrated.
Step 1: Public Parameter Generation / Smart Contract Deployment
Before the genesis block, the creator chooses a security parameter $\lambda$ and runs $\mathsf{pp}\leftarrow\mathsf{SetupCom}(1^{\lambda})$. The parameters $(\lambda,\mathsf{pp})$ are embedded as chain parameters. Once the chain is live, a marketplace smart contract is deployed to record commitments and interactions.
Step 2: Data Validation
Before responding to quotes, a data owner must validate their data with a data authority. The data owner sends $(x,y)$ to the authority for validation. The authority verifies authenticity and ownership and, if successful, computes $\mathtt{com}_{x},r_{x}\leftarrow\mathsf{Commit}(\mathsf{pp},x)$ and $\mathtt{com}_{y},r_{y}\leftarrow\mathsf{Commit}(\mathsf{pp},y)$, then submits $(\mathtt{com}_{x},\mathtt{com}_{y})$ to the marketplace contract from the authority’s contract address (serving as a signature). The authority notifies the data owner of the outcome and, if successful, returns $(r_{x},r_{y})$.
Step 3: Model Validation
Model owners prepare a model suitable for evaluation (including preprocessing such as distillation and splitting) and submit it to a model authority for validation. The authority verifies the reconstruction quality at the BBCC split. If successful, it computes $\mathtt{com}_{\theta_{s}},r_{s}\leftarrow\mathsf{Commit}(\mathsf{pp},\theta_{s})$ for $s\in\{A,B,C\}$ and submits
$(\mathtt{com}_{\theta_{A}},\mathtt{com}_{\theta_{B}},\mathtt{com}_{\theta_{C}})$ on-chain from the model authority’s contract address. The authority notifies the model owner and, if successful, returns $(r_{A},r_{B},r_{C})$.
Step 4: Request Data Quotes
A model owner requests quotes by calling the marketplace contract, which emits an event that contributors can monitor. The request includes data requirements, model commitments $(\mathtt{com}_{A},\mathtt{com}_{B},\mathtt{com}_{C})$, and an escrow that is forfeited if the model owner deviates from the protocol.
Step 5: Submit Data Quotes
Data contributors respond by submitting their commitments $(\mathtt{com}_{x},\mathtt{com}_{y})$ and an escrow to the contract. Submissions must match the authority-published commitments from Step 2; otherwise, the quote is automatically rejected.
Step 6: Data Evaluation Protocol
For each accepted quote, the data owner and model owner run PrivaDE, with two blockchain-oriented refinements:
  1. (1)
    MPC fairness. In $\Pi_{\mathsf{SubScore}}$, the final protocol message and output can be escrow-enforced: the last-sender must pre-commit (e.g., via an on-chain encrypted blob or hash-locked reference) to the final message/output by a deadline. Failure to comply triggers escrow slashing and compensates the counterparty, preventing an unfair abort after learning $\phi$.
  2. (2)
    Payment for ZKP generation. In $\mathsf{CP}$ and $\mathsf{CnCZK}$, verifiers may over-request proofs to raise the security if there is no cost attached. The contract can require proportional payment from the verifier before proof generation, aligning incentives and reducing prover load.
Step 7: Data Transaction
After scoring (potentially across multiple contributors), the model owner proceeds with transactions to selected data owners. Transactions can use hash-locked exchanges (Poon and Dryja, 2016):
  1. (1)
    The data owner uploads an encrypted dataset to IPFS, submits $H(k)$ on-chain, and privately sends the CID and a ZKP attesting that the ciphertext decrypts under $k$;
  2. (2)
    The model owner retrieves the CID, verifies the ZKP, and escrows payment to the contract;
  3. (3)
    The data owner reveals $k$ on-chain; the contract checks $H(k)$, releases escrow to the data owner, and the model owner decrypts the data—completing the transaction.
7. Evaluation
To assess feasibility, we benchmark the secure protocol with respect to runtime, memory usage, communication cost, and related metrics. Scalability is analyzed by varying both model size and dataset complexity. We further compare our active data selection strategy against entropy-based sampling and the Core-Set method (Sener and Savarese, 2018), evaluating its effectiveness in selecting batches that most improve model performance.
Our open-source implementation is publicly available. We instantiate SMPC using the SPDZ2k protocol within MP-SPDZ (Keller, 2020), employ EZKL (zkonduit, 2023) for ZKP-based inference, and implement the challenge protocol in Circom (Bellés-Muñoz et al., 2023).
We validate on MNIST (LeCun and Cortes, 2010), CIFAR-10 (Krizhevsky, 2009), and CIFAR-100, testing three representative models: LeNet (61,706 parameters), ResNet-20 (272,474 parameters), and VGG-8 (4,506,276 parameters).
7.1. Practicality of PrivaDE
We implement PrivaDE in a two‐party simulation. Each party executes as a separate process on a single host (32‐core Intel Core i9, 128 GB RAM), communicating over distinct TCP ports. The loss function is always cross‐entropy $\mathcal{L}_{\mathrm{CE}}$; uncertainty is computed as average prediction entropy; diversity is computed as average feature‐wise standard deviation. The representative set is selected via the kk‐center greedy algorithm.
An overview of the experimental setup is shown in Table 1. For each dataset we employed models of increasing size, then applied distillation to reduce them. We did not tune model accuracy deliberately, reflecting real-world scenarios where data shortages motivate evaluation of new contributions.
Table 2 reports results of the $\mathsf{CP}$ protocol for representative set verification. Setup time refers to SRS generation, key generation, and program compilation, while online time includes both proof generation and verification. Results show stable online performance across experiments, as $\mathsf{CP}$ runtime is independent of model size; all online times remain below 60 seconds.
The model split procedure (A–B–C) is assumed precomputed and validated (see App. D). Table 3 presents performance of part A inference within malicious 2PC. Online timings are consistently low (sub-20 seconds), with the offline phase dominating due to heavy communication overhead. Since offline time consists primarily of input-independent randomness generation, it can be completed prior to execution.
Inference for parts B and C, executed locally but verified via $\mathsf{CnCZK}$ (Tables 3 and 4), show modest resource requirements: both setup and online costs are low. This demonstrates that cut-and-choose significantly reduces time, memory, and communication overhead while still providing reasonable assurance.
Finally, Table 4 shows the timings of 2PC data scoring. Runtimes are stable, though CIFAR-100 incurs higher overhead due to its larger output dimension.
Overall, Table 5 summarizes the end-to-end runtime of the full protocol. While PrivaDE still requires nontrivial resources, it is feasible: smaller models require only about one minute of online time and less than 15 minutes end-to-end. Further optimizations—such as reducing the size of block A or lowering the number of $\mathsf{CP}$ and $\mathsf{CnCZK}$ checks—can trade off security assurance for reduced runtime, depending on application needs.
Table 1. Overview of datasets and model variants tested in section 7.1. Models are tested with increasing size; with distillation that brings 60-90% reduction in model sizes.
Dataset Original Model Distilled Model Bob Dataset Size Original Model Size (# Params) Distilled Model Size (# Params) Reduction (%) Teacher Accuracy Student Accuracy
MNIST LeNet-5 LeNetXS 1000 61,706 3,968 93.57 0.91 0.91
CIFAR-10 ResNet20 LeNet-5 1000 272,474 83,126 69.49 0.37 0.27
CIFAR-100 VGG8 5-Layer CNN 1000 4,506,276 1,727,588 61.66 0.1 0.05
Table 2. Overview of parameters and runtime performances of the $\mathsf{CP}$ challenge protocol. Results are stable across different experiments as the protocol is independent of model sizes.
Dataset Representative Set Size Data Dimension # Challenges $\mathsf{CP}$ Setup Time (s) $\mathsf{CP}$ Online Time (s) Memory Usage (MB) Communication Overhead (MB)
MNIST 50 35 20 456 28.0 2,579 0.0153
CIFAR-10 50 50 20 435 50.1 2,505 0.0154
CIFAR-100 50 50 20 435 56.1 2,501 0.0154
Table 3. Model A and Model B inference performance across experiments. 2PC inference of model A incurs significant communication overhead but online runtimes are small. ZKP verification required for model B inference are all within practical runtimes.
Model A (2PC) Model B ($\mathsf{CnCZK}$)
Dataset Online (s) Offline (s) Comm. (MB) % of Computations Verified Layers Sampled Data points Sampled Setup Time (s) Online Time (s) Memory Usage (MB) Comm. (MB)
MNIST 4.43 193 402,343 30 1/2 30/50 2.12 2.29 4,130 0.353
CIFAR-10 5.14 193 346,774 25.7 3/7 30/50 5.75 6.63 5,508 1.25
CIFAR-100 20.2 761 1,355,480 30 1/2 30/50 19 15 10,429 6.1
Table 4. Model C and 2PC Scoring performance across experiments. Model C inference verification scaled with the number of layers sampled, while 2PC scoring scaled with the number of classes.
Model C ($\mathsf{CnCZK}$) 2PC Data Scoring
Dataset % Comp. Verified Layers Sampled Data points Sampled Setup Time (s) Online Time(s) Memory Usage (MB) Comm. (MB) Offline Time (s) Online Time (s) Comm. (MB)
MNIST 0.257 3/7 30/50 9.56 9.84 4,312 1.05 90.6 0.45 33,825
CIFAR-10 0.2 1/3 30/50 3.42 3.57 6,314 0.362 90.9 0.44 33,825
CIFAR-100 0.3 11/22 30/50 657 779 14,962 60.9 618 2.94 230,798
Table 5. Overall running times per dataset. Results show a satisfactory online runtime.
Dataset Online (s) Total (s)
MNIST 45.0 778
CIFAR-10 65.9 794
CIFAR-100 873.0 3,360
7.2. Robustness of the scoring algorithm
The multi‐point scoring algorithm (Algorithm 1) is evaluated against entropy‐based sampling (Settles and Craven, 2008) and core‐set sampling (Sener and Savarese, 2018) on MNIST, CIFAR-10 and CIFAR-100. After normalization and Gaussian blurring, each dataset is split into a pre‐training set (Alice’s initial data), a held‐out test set, and $N$ contributing sets with skewed class distributions to simulate heterogeneity.
Using LeNet for MNIST, ResNet20 for CIFAR-10 and VGG-8 for CIFAR-100 (same setup as above), our framework comprises: representative selection $\mathcal{S}$ via K‐Center‐Greedy (Sener and Savarese, 2018); loss $\ell=\mathcal{L}_{\mathrm{CE}}$; uncertainty $\mathcal{U}$ as Shannon entropy; diversity
$\mathcal{D}(D_{B}^{x},D_{R}^{x})=\max_{a\in D_{B}^{x}\setminus D_{R}^{x}}\min_{b\in D_{R}^{x}}\|a-b\|$; and scoring $f(\ell,u,d)=\alpha_{1}\ell+\alpha_{2}u+\alpha_{3}d$ with $\alpha_{1}=0.2$, $\alpha_{2}=0.1$, $\alpha_{3}=0.7$ via grid search.
We simulate sequential acquisition: at each iteration, Alice scores each contributing set, acquires the top batch via our protocol, augments her training data, retrains, and records test accuracy until all $N$ batches are processed. For baselines, entropy sampling uses mean batch entropy, while core‐set sampling selects ten cluster centers (K‐Center‐Greedy) and scores batches by the maximum sample‐to‐center distance. Accuracy curves are shown in Figure 10.
Figure 10. Robustness of our multi-point scoring algorithm vs. other active learning methods. The results show our method is comparable to core-set and outperforms random and entropy sampling.
Analysis. As Figure 10 shows, our method consistently outperforms random and entropy sampling. On MNIST (Figure 10(a)), it yields the largest early gains; on CIFAR-10 (Figure 10(b)), it matches core‐set sampling. These results highlight that, under batch constraints, diversity drives early improvements more than uncertainty, and integrating loss, uncertainty, and diversity produces a robust selection criterion across architectures and datasets.
Effect of using distilled models. We further assess the impact of using a distilled model in place of the original teacher model by measuring the Spearman correlation between the scores produced by both. Using the same setup as in the feasibility tests, Table 6 reports the results: all teacher–student pairs exhibit high correlation. The only exception is a reduced correlation in the CIFAR-100 setting, which we attribute to the higher class uncertainty and lower baseline accuracy of the original model.
Table 6. Correlation between teacher and student produced scores across 5 runs (mean $\pm$ SD). Results show that all models generally have high correlation, with slightly lower correlation seen in CIFAR-100, possibly attributing to lower accuracy of the initial model.
Dataset Teacher Student Mean $\pm$ SD
MNIST LeNet5 LeNetXS 0.8742 $\pm$ 0.0592
CIFAR-10 resnet20 LeNet5 0.8478 $\pm$ 0.1199
CIFAR-100 VGG8 5-layer CNN 0.6285 $\pm$ 0.1158
8. Related Works
Privacy-preserving Data Evaluation. Prior work has explored privacy-preserving data valuation and selection.
Song et al. (Song et al., 2021) propose a data evaluation protocol based on uncertainty sampling, later refined by Qian et al. (Qian et al., 2024). Their main advance is an inner-product-based functional encryption scheme that accelerates inference. However, both works adopt the semi-honest adversarial model, which is insufficient for blockchain-based deployments where no trust is assumed between parties.
Tian et al. (Tian et al., 2022) take a different approach by estimating Shapley values via a learned Data Utility Model and then executing custom MPC to perform inference and approximate Shapley scores. This method, however, requires sharing portions of the data owner’s samples and additional labeling to train the utility model; moreover, the end-to-end MPC cost of full-model inference remains high (e.g., $\sim$16 hours for 2,000 CIFAR-10 images in their evaluation), which is ill-suited to marketplaces that demand quick, low-effort data assessment and purchase.
Data pricing. Data pricing is usually concerned with setting a price for data in a standard currency. The research in this area adapts economic principles of pricing and markets to data (Zhang et al., 2023), and is concerned with the data-specific challenges such as duplication and arbitrage (Chen et al., 2019; Ohrimenko et al., 2019). These works take the perspective of the broker and consider the seller and buyer’s market separately. The actual utility of the data to the consumer is assumed to be determined by simple attributes (Koutris et al., 2015).
Game theoretic fair data valuation. This topic emerged with the popularity of machine learning and demand for data (Jia et al., 2019; Ghorbani et al., 2020; Ghorbani and Zou, 2019), and can be seen as an aspect of explainable machine learning (Rozemberczki et al., 2022). The aim of game theoretic data valuation is to determine fair value for data points (or subsets) in a multiple data contributor setting. These works make use of Shapley Value (Rozemberczki et al., 2022) with its Fairness properties to ensure fair valuation for all. Crucially, in these problems the fairness consideration is between different data contributors, as Shapley Value is a mechanism for determining fair share between multiple players. Several associated approaches exist for evaluating data, such as leave-one-out testing (Cook, 1977) and influence function estimation (Koh and Liang, 2017; Datta et al., 2015; Pruthi et al., 2020), with similar objectives. Shapley value and similar techniques tend to be computationally expensive – repeatedly training and testing models on different subsets of data – and require a trusted party to carry out the computation.
Active learning. We borrowed several ideas from this area, such as the use of uncertainty and diversity (Settles, 2009; Dasgupta, 2011). Methods that use both are called hybrid methods. There are algorithms incorporating different scoring values non-linearly (Ash et al., 2020; Yang et al., 2017; Zhdanov, 2019) and linearly (Cardoso et al., 2017). In the standard active learning paradigm, loss is not a decision criterion, because loss computation requires the ground truth labels, which are assumed to be expensive, and the algorithms aim to minimize the use of labels. In (Yoo and Kweon, 2019), a loss predictor is trained simultaneously with the model to inform the active selection process.
Zero knowledge proofs in Machine learning. Researches in this area have been mainly focused on generalizability and efficiency. For generalizability, there are works mapping machine learning operations into suitable zero-knowledge circuit, and translation of non-linear and floating point operations into suitable approximate encryption-friendly circuits (e.g. (Ghodsi et al., 2017)). Other researches in this area have approached the efficiency problem by making use of specific features of models and training algorithms (Angel et al., 2022). However, current zero knowledge proof systems are considered useful mainly for model inference and they are not efficiently supporting the training process. Commonly they affect the model’s accuracy or can be too slow for training even moderate sized models.
Our setting of data scoring using zero knowledge proofs creates some unique challenges and possibilities. It allows the use of labels and loss in ways that are not possible in standard active learning, and ensures the transmission of enough information to make decision without revealing secret data.
Secure MPC Inference Methods. Although zero‐knowledge proofs (ZKPs) protect model weights and guarantee correctness of inference, they do not by themselves ensure privacy of the query input $x$. An alternative is to employ maliciously‐secure multiparty computation (MPC) under a dishonest–majority model, which provides both confidentiality and integrity in a two‐party setting. The most prominent example here is MD‐ML (Yuan et al., 2024), which optimizes cryptographic primitives to minimize online inference latency. While MD‐ML achieves low absolute latency on small models, its performance degrades sharply with complexity––incurring roughly a 40× slowdown from a Linear SVM to LeNet.
Many other maliciously‐secure inference frameworks assume an honest–majority among three or more parties, which is incompatible with two‐party scenarios. Representative systems in this category include Helix (Zhang et al., 2025b), Blaze (Patra and Suresh, 2020), and SWIFT (Koti et al., 2021). Because these require a strict majority of honest participants, their security and performance guarantees do not directly translate to our setting.
9. Conclusion
Large-scale AI models and training data remain concentrated among major corporations and institutions. Decentralized data marketplaces aim to lower this barrier by establishing trust and incentivizing data sharing, thereby enabling individual data owners and model builders to collaboratively develop advanced models.
This paper introduced PrivaDE, a practical and efficient protocol for privacy-preserving data evaluation in decentralized marketplaces. PrivaDE operates in the malicious-adversary setting with explicitly bounded leakage, combining representative-set selection, verified split inference, and secure scoring to deliver actionable utility signals. Our experiments demonstrate its feasibility and efficiency.
Future directions include optimizing cryptographic subroutines for lower latency and communication, improving scalability to ever-larger datasets and models, and building fully decentralized (blockchain-based) deployments that integrate data validation, payments, fair MPC, and model training into a cohesive end-to-end workflow.
Appendix A Proofs of Theorems
A.1. Proof of Theorem 4.2
In the worst case, for each corrupted point $i$, the adversary modifies only a single value—either one activation $a_{l,i}$ among the L+1L\!+\!1 activations ($l\in\{0,\dots,L\}$) or one layer weight $\theta_{j}$ among the $L$ weights. Because altering activations can be caught with slightly higher probability (if $l\in\{1,\dots,L-1\}$, detection occurs when either layer $l$ or l+1l\!+\!1 is checked), we analyze the stronger adversary that alters only a weight.
Hence, the detection probability for a single corrupted point, conditioned on that point being audited, is bounded by
$p_{\text{single}}\;\geq\;\frac{s}{L}.$
If $k$ corrupted points are audited, the probability of detecting at least one of them is
$1-\Pr[\text{no corrupted point detected}]\;=\;1-(1-p_{\text{single}})^{k}.$
Let $K$ be the number of corrupted points drawn into the audit set $S$; then $K$ follows a hypergeometric distribution with
$\Pr[K=k]\;=\;\frac{\binom{N\rho}{k}\binom{N(1-\rho)}{\,m-k\,}}{\binom{N}{m}},\qquad k=0,1,\dots,\min(N\rho,m).$
Therefore,
$\Pr[\textsf{detect}]$ $=\sum_{k=0}^{\min(N\rho,\,m)}\Pr[K=k]\cdot\bigl[1-(1-p_{\text{single}})^{k}\bigr]$
$\geq\sum_{k=0}^{\min(N\rho,\,m)}\frac{\binom{N\rho}{k}\binom{N(1-\rho)}{\,m-k\,}}{\binom{N}{m}}\left[1-\left(\tfrac{L-s}{L}\right)^{k}\right],$
which proves the claim.
Appendix B Security Definitions
B.1. Commitment Schemes
Definition B.1 (Commitment Schemes).
A commitment scheme is a tuple $\Gamma=(\mathsf{Setup},\mathsf{Commit},\mathsf{Open})$ of $\mathsf{PPT}$ algorithms where:
  • $\mathsf{Setup}(1^{\lambda})\rightarrow\mathsf{pp}$ takes security parameter $\lambda$ and generates public parameters pp;
  • $\mathsf{Commit}(\mathsf{pp};m)\rightarrow(\mathtt{com},r)$ takes a secret message $m$, outputs a public commitment $Com$ and a randomness $r$ used in the commitment.
  • $\mathsf{Open}(\mathsf{pp},\mathtt{com};m,r)\rightarrow b\in\{0,1\}$ verifies the opening of the commitment $C$ to the message $m$ with the randomness $r$.
$\Gamma$ has the following properties:
  • Binding. For all $\mathsf{PPT}$ adversaries $\mathcal{A}$, it holds that:
    $\Pr\left[b_{0}=b_{1}\neq 0\land m_{0}\neq m_{1}:\begin{array}[]{l}\mathsf{pp}\leftarrow\mathsf{Setup}(1^{\lambda})\\ (\mathtt{com},m_{0},m_{1},r_{0},r_{1})\leftarrow\mathcal{A}(\mathsf{pp})\\ b_{0}\leftarrow\mathsf{Open}(\mathsf{pp},\mathtt{com},m_{0},r_{0})\\ b_{1}\leftarrow\mathsf{Open}(\mathsf{pp},\mathtt{com},m_{1},r_{1})\end{array}\right]\leq\nu(\lambda)$
  • Hiding. For all $\mathsf{PPT}$ adversaries $\mathcal{A}$, it holds that:
    $\left|\Pr\left[b_{0}=b^{\prime}:\begin{array}[]{l}\text{pp}\leftarrow\mathsf{Setup}(1^{\lambda})\\ (m_{0},m_{1},\text{st})\leftarrow\mathcal{A}(\text{pp})\\ b\overset{\mathdollar}{\leftarrow}\{0,1\}\\ (C_{b},r_{b})\leftarrow\mathsf{Commit}(\text{pp},m_{b})\\ b^{\prime}\leftarrow\mathcal{A}(\text{pp},\text{st},C_{b})\end{array}\right]-\frac{1}{2}\right|=\nu(\lambda)$
B.2. Merkle Tree
Definition B.2 (Merkle Tree Commitments).
Fix a collision-resistant hash function family $\mathcal{H}=\{H_{\text{pp}}:\{0,1\}^{*}\to\{0,1\}^{\kappa}\}$ with public parameter pp and domain-separation tags $\mathsf{leaf},\mathsf{node}\in\{0,1\}^{*}$. A (binary) Merkle tree commitment scheme is a tuple
$\Delta=(\mathsf{mtSetup},\ \mathsf{mtCommit},\ \mathsf{mtOpen},\ \mathsf{mtVerify})$
of $\mathsf{PPT}$ algorithms where:
  • $\mathsf{mtSetup}(1^{\lambda})\rightarrow\text{pp}$: samples pp for $H_{\text{pp}}\leftarrow\mathcal{H}$
  • $\mathsf{mtCommit}(\text{pp};\ m_{1},\dots,m_{n})\rightarrow(\mathtt{com},\mathsf{\vphantom{p}st})$: on inputs $m_{1},\dots,m_{n}$ (the “leaves”), compute for each $i\in[n]$ a leaf hash
    $\ell_{i}\;=\;H_{\text{pp}}\bigl(\mathsf{leaf}\,\|\,\mathsf{enc}(i)\,\|\,m_{i}\bigr)$
    , then iteratively compute internal nodes
    $v\;=\;H_{\text{pp}}\bigl(\mathsf{node}\,\|\,v_{L}\,\|\,v_{R}\bigr)$
    up to the root $v_{\mathsf{root}}$. Output the commitment $\mathtt{com}\leftarrow v_{\mathsf{root}}$ and opening state $\mathsf{\vphantom{p}st}$ containing the tree (or any data sufficient to derive authentication paths).
  • $\mathsf{mtOpen}(\text{pp},\mathsf{\vphantom{p}st};\ i)\rightarrow\pi_{i}$: outputs a Merkle opening for position $i\in[n]$, namely the authentication path
    $\pi_{i}\;=\;\bigl((s_{1},b_{1}),\dots,(s_{h},b_{h})\bigr),$
    where $s_{j}$ is the sibling hash at level $j$ and $b_{j}\in\{L,R\}$ indicates whether the running value is a left or right child; $h=\lceil\log_{2}n\rceil$.
  • $\mathsf{mtVerify}(\text{pp},\mathtt{com};\ i,m_{i},\pi_{i})\rightarrow b\in\{0,1\}$: recompute
    $x_{0}\leftarrow H_{\text{pp}}\bigl(\mathsf{leaf}\,\|\,\mathsf{enc}(i)\,\|\,m_{i}\bigr),\ \ x_{j}\leftarrow\begin{cases}H_{\text{pp}}\bigl(\mathsf{node}\,\|\,x_{j-1}\,\|\,s_{j}\bigr)&\text{if }b_{j}=L,\\ H_{\text{pp}}\bigl(\mathsf{node}\,\|\,s_{j}\,\|\,x_{j-1}\bigr)&\text{if }b_{j}=R,\end{cases}$
    and output $b=1$ iff $x_{h}=\mathtt{com}$.
Let $\Delta$ be as in Definition B.2. Then:
  • Completeness. For all $n$ and $\vec{m}\!\in\!(\{0,1\}^{*})^{n}$, if $\text{pp}\!\leftarrow\!\mathsf{mtSetup}(1^{\lambda})$, $(\mathtt{com},\mathsf{\vphantom{p}st})\!\leftarrow\!\mathsf{mtCommit}(\text{pp},\vec{m})$, and $\pi_{i}\!\leftarrow\!\mathsf{mtOpen}(\text{pp},\mathsf{\vphantom{p}st};i)$, then $\mathsf{mtVerify}(\text{pp},\mathtt{com};i,m_{i},\pi_{i})=1$ for every $i\in[n]$.
  • Position-Binding. For all $\mathsf{PPT}$ adversaries $\mathcal{A}$, the probability that $\mathsf{mtVerify}(\text{pp},\mathtt{com};i,m_{0},\pi_{0})=\allowbreak\mathsf{mtVerify}(\text{pp},\mathtt{com};i,m_{1},\pi_{1})\allowbreak=1$ with $m_{0}\neq m_{1}$ is at most $\nu(\lambda)$, given $\text{pp}\!\leftarrow\!\mathsf{mtSetup}(1^{\lambda})$ and $(\mathtt{com},i,m_{0},\pi_{0},m_{1},\pi_{1})\!\leftarrow\!\mathcal{A}(\text{pp})$. (By collision resistance of $H$.)
B.3. Zero-Knowledge Proofs
Definition B.3 (Non-Interactive Zero-Knowledge proof-of-knowledge Systems).
A non-interactive zero-knowledge proof-of-knowledge system (NIZKPoK) for an NP-language $L$ with the corresponding relation $\mathcal{R}_{L}$ is a non-interactive protocol $\Pi=(\mathsf{Setup},\mathcal{P},\mathcal{V})$, where:
  • $\mathsf{Setup}(1^{\lambda})\rightarrow\mathtt{CRS}$ takes as the input a security parameter $\lambda$. It outputs a common reference string $\mathtt{CRS}$.
  • $\mathcal{P}(\mathtt{CRS},x,w)\rightarrow\pi$ takes as the input $\mathtt{CRS}$, the statement $x$ and the witness $w$, s.t. $(x,w)\in\mathcal{R}_{L}$. It outputs the proof $\pi$.
  • $\mathcal{V}(\mathtt{CRS},x,\pi)\rightarrow b\in\{0,1\}$ takes as the input $\mathtt{CRS}$, $x$ and $\pi$. It outputs 1 to accept and 0 to reject.
$\Pi$ has the following properties:
  • Completeness. For all $\lambda\in\mathbb{N}$, and all $(x,w)\in\mathcal{R}_{L}$, it holds that:
[…]
  • Soundness. For all $\mathsf{PPT}$ provers $\mathcal{P}^{\star}$, s.t. for all $\lambda\in\mathbb{N}$, and all $x\notin L$, it holds that:
[…]
  • Zero knowledge. There exists a $\mathsf{PPT}$ simulator $\mathsf{Sim}$ such that for every $(x,w)\in\mathcal{R}_{L}$, the distribution ensembles $\{(\mathtt{CRS},\pi):\mathtt{CRS}\overset{\mathdollar}{\leftarrow}\mathsf{Setup}(1^{|x|},1^{\lambda});$ $\pi\overset{\mathdollar}{\leftarrow}\mathcal{P}(\mathtt{CRS},x,w)\}_{\lambda\in\mathbb{N}}$ and $\{\mathsf{Sim}(1^{\lambda},x)\}_{\lambda\in\mathbb{N}}$ are computationally indistinguishable.
  • Proof of Knowledge. For all $\mathsf{PPT}$ provers $\mathcal{P}$, there exists an extractor $\mathcal{E}$ such that
[…]
B.4. Multi-Party Computation
Here, we provide a formal definition of secure multiparty computation, with emphasis on the two-party scenario.
A multiparty protocol is cast by specifying a random process that maps tuples of inputs to tuples of outputs (one for each party). We refer to such a process as a functionality. The security of a protocol is defined with respect to a functionality $f$. In particular, let $n$ denote the number of parties. A (non-reactive) $n$-party functionality $f$ is a (possibly randomized) mapping of $n$ inputs to $n$ outputs. A multiparty protocol with security parameter $\lambda$ for computing the functionality $f$ is a protocol running in time $\mathsf{poly}(\lambda)$ and satisfying the following correctness requirement: if parties $P_{1},\dots,P_{n}$ with inputs $x_{1},\dots,x_{n}$ respectively, all run an honest execution of the protocol, then the joint distribution of the outputs $y_{1},\dots,y_{n}$ of the parties is statistically close to $f(x_{1},\dots,x_{n})$.
The security of a protocol (w.r.t. a functionality $f$) is defined by comparing the real-world execution of the protocol with an ideal-world evaluation of $f$ by a trusted party. More concretely, it is required that for every adversary $\mathcal{A}$, which attacks the real-world execution of the protocol, there exists an adversary $\mathcal{S}$, also referred to as the simulator, which can achieve the same effect in the ideal-world execution. We denote $\vec{x}=(x_{1},\dots,x_{n})$.
Malicious MPC. In the real execution for the malicious assumption, the $n$-party protocol $\Pi$ for computing $f$ is executed in the presence of an adversary $\mathcal{A}$. The honest parties follow the instructions of $\Pi$. The adversary $\mathcal{A}$ takes as input the security parameter $\lambda$, the set $I\subset[n]$ of corrupted parties, the inputs of the corrupted parties, and an auxiliary input $z$. $\mathcal{A}$ sends all messages in place of corrupted parties and may follow an arbitrary polynomial-time strategy.
[…]
An ideal execution for a function $f$ proceeds as follows:
  • Send inputs to the trusted party: As before, the parties send their inputs to the trusted party, and we let $x^{\prime}_{i}$ denote the value sent by $P_{i}$.
  • Trusted party sends output to the adversary: The trusted party computes $f(x^{\prime}_{1},\dots,x^{\prime}_{n})=(y_{1},\dots,y_{n})$ and sends $\{y_{i}\}_{i\in I}$ to the adversary.
  • Adversary instructs trusted party to abort or continue: This is formalized by having the adversary send either a continue or abort message to the trusted party. (A semi-honest adversary never aborts.) In the former case, the trusted party sends to each uncorrupted party $P_{i}$ its output value $y_{i}$. In the latter case, the trusted party sends the special symbol $\bot$ to each uncorrupted party.
  • Outputs: $\mathcal{S}$ outputs an arbitrary function of its view, and the honest parties output the values obtained from the trusted party.
In the case of secure two-party computation, where an adversary corrupts only one of the two parties, the above behavior of an ideal-world trusted party is captured by functionality $\mathcal{F}_{\sf 2PC}$, described in Figure 11 below.
Parameters. Description of function $f$. Let $\mathcal{H}\in\{1,2\}$ and $\mathcal{M}\in\{1,2\}\setminus\{\mathcal{H}\}$ denote the index of honest party and corrupt party, respectively. (1) $\mathcal{F}_{\sf 2PC}$ receives input $x_{\mathcal{H}}$ from honest party $P_{\mathcal{H}}$ and input $x_{\mathcal{M}}$ of the corrupt party from simulator $\mathcal{S}$. (2) $\mathcal{F}_{\sf 2PC}$ computes $y\leftarrow f(x_{1},x_{2})$ and sends $y$ to $\mathcal{S}$. (3) If $\mathcal{F}_{\sf 2PC}$ receives $\mathsf{abort}$ from $\mathcal{S}$, it sends $\mathsf{abort}$ to honest party $P_{\mathcal{H}}$. Otherwise, $\mathcal{F}_{\sf 2PC}$ sends $y$ to honest party $P_{\mathcal{H}}$.
Figure 11. Functionality $\mathcal{F}_{\sf 2PC}$ for maliciously secure 2PC.
As before, the interaction of $\mathcal{S}$ with the trusted party defines a random variable $\mathsf{IDEAL}_{f,\mathcal{S}(z),I}(\lambda,\vec{x})$. Having defined the real and the ideal world executions, we can now proceed to define the security notion.
Definition B.4.
Let $\lambda$ be the security parameter, $f$ an $n$-party randomized functionality, and $\Pi$ an $n$-party protocol for $n\in\mathbb{N}$. We say that $\Pi$ $t$-securely realizes $f$ in the presence of malicious adversaries if for every PPT adversary $\mathcal{A}$ there exists a PPT adversary $\mathcal{S}$ such that, for any $I\subset[n]$ with $|I|\leq t$, the following quantity is negligible in $\lambda$:
$|\Pr[_{\Pi,\mathcal{A}(z),I}(\lambda,\vec{x})=1]-\Pr[\mathsf{IDEAL}_{f,\mathcal{S}(z),I}(\lambda,\vec{x})=1]|,$
where $\vec{x}=\{x_{i}\}_{I\in[n]}\in\{0,1\}^{\ast}$ and $z\in\{0,1\}^{\ast}$.
Note that for two-party computation, $n=2$ and $t=1$ in Definition B.4.
Appendix C Security Proofs
C.1. Proof of Thm. 5.1
Let $\mathcal{A}$ be a PPT adversary. We first consider the case in which $P_{2}$ (Bob) is corrupted (i.e., $\mathcal{A}$ controls $P_{2}$). The simulator $\mathcal{S}_{\mathsf{Bob}}$ works as follows:
  1. (1)
    $\mathcal{S}_{\mathsf{Bob}}$ computes $(\mathtt{com}^{A}_{0},r^{A}_{0})\leftarrow\mathsf{Commit}(\mathsf{pp},0)$ and $(\mathtt{com}^{C}_{0},r^{C}_{0})\leftarrow\mathsf{Commit}(\mathsf{pp},0)$, then invokes $\mathcal{A}$ with inputs $\mathtt{com}^{A}_{0},\mathtt{com}^{C}_{0}$.
  2. (2)
    $\mathcal{A}$ replies with $\mathtt{com}_{x_{i}},\mathtt{com}_{y_{i}}$ for $i\in[n]$, and the representative set indices $I_{R}$.
  3. (3)
    $\mathcal{S}_{\mathsf{Bob}}$ sends $I_{R}$ to the trusted party to obtain $I$, and forwards $I$ to $\mathcal{A}$.
  4. (4)
    $\mathcal{S}_{\mathsf{Bob}}$ receives the list of ZKPs $\pi_{i}$ from $\mathcal{A}$. For each $\pi_{i}$, it verifies the proof and runs the extractor $(j,x_{i},r_{i},x_{j},r_{j})\leftarrow\mathcal{E}((\mathsf{pp},d,i,\{\mathtt{com}_{x_{i}}\},I_{R}),\mathcal{A})$.
  5. (5)
    $\mathcal{S}_{\mathsf{Bob}}$ sends $D^{x}_{R|I}=\{x_{j}\}_{j=1}^{m}$ and $D^{x}_{B|I}=\{x_{i}\}_{i=1}^{m}$ to the trusted party. If the trusted party replies with $\mathsf{abort}$, $\mathcal{S}_{\mathsf{Bob}}$ aborts.
  6. (6)
    $\mathcal{S}_{\mathsf{Bob}}$ invokes $\mathcal{S}^{\mathsf{Inf}}_{\mathsf{Bob}}$ for subprotocol $\Pi_{\mathsf{Inference}}$ to obtain $D^{x}_{R},\{r_{i}\}_{i}$. It then sends $D^{x}_{R}$ to the trusted party, receives activations $\overrightarrow{a_{A}}$, and computes commitments $\mathtt{com}_{a_{i}},r_{a_{i}}\leftarrow\mathsf{Commit}(\mathsf{pp},a_{i})$.
  7. (7)
    It checks each commitment $\mathsf{Open}(\mathsf{pp},\mathtt{com}_{i},x_{i},r_{i})$ for $i\in I_{R}$. If any are invalid, it sends $\mathsf{abort}$ to $\mathcal{S}^{\mathsf{Inf}}_{\mathsf{Bob}}$. Otherwise, it sends $(\overrightarrow{a},\{\mathtt{com}_{a_{i}},r_{a_{i}}\}_{i})$ to $\mathcal{S}^{\mathsf{Inf}}_{\mathsf{Bob}}$ and forwards its output to $\mathcal{A}$.
  8. (8)
    $\mathcal{S}_{\mathsf{Bob}}$ receives $(\overrightarrow{a_{B}},\pi_{B})$ from $\mathcal{A}$ and $(\theta_{B},\overrightarrow{a_{B}}^{\prime})$ from the trusted party. It verifies $\pi_{B}$ (for $L_{\textsf{ML-Inf}}$ with public weights, hidden inputs) using $(\mathsf{pp},\theta_{B},\{\mathtt{com}_{a_{i}}\},\overrightarrow{a_{B}})$ and checks $\overrightarrow{a_{B}}=\overrightarrow{a_{B}}^{\prime}$. If verification fails, it aborts.
  9. (9)
    $\mathcal{S}_{\mathsf{Bob}}$ creates $k$ dummy commitments $(\mathtt{com}_{y^{\prime}_{i}},r_{y^{\prime}_{i}})\leftarrow\mathsf{Commit}(\mathsf{pp},0)$ for $i\in[k]$. It then invokes $\mathcal{S}_{\mathsf{ZK}}$ for $L_{\textsf{ML-Inf}}$ (hidden weights, public inputs) to generate $\pi\leftarrow\mathcal{S}_{\mathsf{ZK}}(\mathsf{pp},\overrightarrow{a_{B}},\mathtt{com}^{C}_{0},\{\mathtt{com}_{y^{\prime}_{i}}\}_{i})$. It sends $\pi,\mathtt{com}_{y^{\prime}_{1}},\ldots,\mathtt{com}_{y^{\prime}_{k}}$ to $\mathcal{A}$. If $\mathcal{A}$ aborts, it aborts.
  10. (10)
    $\mathcal{S}_{\mathsf{Bob}}$ invokes $\mathcal{A}$ to obtain $(D_{B}^{x},y_{1},\ldots,y_{k},r_{y_{1}},\ldots,r_{y_{k}},\mathtt{com}_{y^{\prime}_{1}},\ldots,\mathtt{com}_{y^{\prime}_{k}})$ for scoring.
  11. (11)
    $\mathcal{S}_{\mathsf{Bob}}$ sends $(D_{B}^{x},D_{R}^{y})$ to the trusted party and receives score $\phi$.
  12. (12)
    Finally, $\mathcal{S}_{\mathsf{Bob}}$ checks $\mathsf{Open}(\mathsf{pp},\mathtt{com}_{y_{i}},y_{i},r_{y_{i}})$ for each $\mathtt{com}_{y_{i}}$. If any fail, it invokes $\mathcal{S}^{\mathsf{Score}}_{\mathsf{Bob}}$ for $\Pi_{\mathsf{SubScore}}$ with input $\mathsf{abort}$, otherwise with $\phi$. It outputs whatever $\mathcal{S}^{\mathsf{Score}}_{\mathsf{Bob}}$ outputs.
We now define the hybrids:
  • $\mathcal{H}^{b}_{0}$: Real-world execution.
  • $\mathcal{H}^{b}_{1}$: Messages for $\Pi_{\mathsf{SubScore}}$ are computed by $\mathcal{S}^{\mathsf{Score}}_{\mathsf{Bob}}$. Transition by security of $\Pi_{\mathsf{SubScore}}$.
  • $\mathcal{H}^{b}_{2}$: The ZKP $\pi$ for model $C$ inference is simulated by $\mathcal{S}_{\mathsf{ZK}}$. Transition by zero-knowledge.
  • $\mathcal{H}^{b}_{3}$: Model weights $\theta_{A},\theta_{C}$ are replaced by 0; inference computed as $y^{\prime}\leftarrow C_{0}(x)$ and committed. Transition by hiding of commitments.
  • $\mathcal{H}^{b}_{4}$: Inference of C is skipped; commitments of $y^{\prime}$ replaced by commitments of 0. Transition by hiding.
  • $\mathcal{H}^{b}_{5}$: Messages for $\Pi_{\mathsf{Inference}}$ are simulated by $\mathcal{S}^{\mathsf{Inf}}_{\mathsf{Bob}}$. Transition by security of $\Pi_{\mathsf{Inference}}$.
Each transition is justified by standard properties of the underlying protocols. Hence $\mathcal{H}^{b}_{0}\approx\mathcal{H}^{b}_{5}$, and the real and ideal executions are indistinguishable when Bob is corrupted.
Now consider the case where Alice is corrupted (i.e., $\mathcal{A}$ controls $P_{1}$). The simulator $\mathcal{S}_{\mathsf{Alice}}$ works as follows:
  1. (1)
    $\mathcal{S}_{\mathsf{Alice}}$ invokes $\mathcal{A}$, which replies with $\mathtt{com}_{A},\mathtt{com}_{C}$.
  2. (2)
    For each $i\in[n]$, it creates dummy commitments $(\mathtt{com}_{x_{i}},r_{x_{i}})\leftarrow\mathsf{Commit}(\mathsf{pp},0)$ and $(\mathtt{com}_{y_{i}},r_{y_{i}})\leftarrow\mathsf{Commit}(\mathsf{pp},0)$, sending them to $\mathcal{A}$.
  3. (3)
    $\mathcal{S}_{\mathsf{Alice}}$ queries the trusted party for $I_{R}$ and forwards $I_{R}$ to $\mathcal{A}$.
  4. (4)
    $\mathcal{A}$ replies with indices $I$. $\mathcal{S}_{\mathsf{Alice}}$ forwards $I$ to the trusted party. If it replies with $\mathsf{abort}$, $\mathcal{S}_{\mathsf{Alice}}$ aborts.
  5. (5)
    For each $i\in I$, it invokes $\mathcal{S}_{\mathsf{ZK}}$ of $\mathsf{CP}$ to generate $\pi_{i}\leftarrow\mathcal{S}_{\mathsf{ZK}}(\mathsf{pp},d,i,\{com_{x_{j}}\}_{j},I_{R})$ and sends all $\pi_{i}$ to $\mathcal{A}$. If $\mathcal{A}$ aborts, it aborts.
  6. (6)
    $\mathcal{S}_{\mathsf{Alice}}$ invokes $\mathcal{S}^{\mathsf{Inf}}_{\mathsf{Alice}}$ to obtain $(\theta_{A},r_{A})$ and sends $\theta_{A}$ to the trusted party.
  7. (7)
    It checks $\mathsf{Open}(\mathsf{pp},\mathtt{com}_{A},\theta_{A},r_{A})$. If invalid, it sends $\mathsf{abort}$ to $\mathcal{S}^{\mathsf{Inf}}_{\mathsf{Alice}}$. Otherwise, it creates dummy commitments $\mathtt{com}_{a_{i}},r_{a_{i}}\leftarrow\mathsf{Commit}(\mathsf{pp},0)$ and sends $\{\mathtt{com}_{a_{i}}\}$ to $\mathcal{S}^{\mathsf{Inf}}_{\mathsf{Alice}}$, forwarding its output to $\mathcal{A}$.
  8. (8)
    $\mathcal{S}_{\mathsf{Alice}}$ sends $\theta_{B}$ to the trusted party, receives $\overrightarrow{a_{B}}$, and invokes $\mathcal{S}^{\prime}_{\mathsf{ZK}}$ for $L_{\textsf{ML-Inf}}$ (hidden inputs, public weights) to generate $\pi\leftarrow\mathcal{S}^{\prime}_{\mathsf{ZK}}(\mathsf{pp},\theta_{B},\{\mathtt{com}_{a_{i}}\}_{i},\overrightarrow{a_{B}})$. It sends $(\overrightarrow{a_{B}},\pi)$ to $\mathcal{A}$. If $\mathcal{A}$ aborts, it aborts.
  9. (9)
    $\mathcal{S}_{\mathsf{Alice}}$ queries $\mathcal{A}$ for $\pi_{C}$ and commitments $\mathtt{com}_{y^{\prime}_{1}},\ldots,\mathtt{com}_{y^{\prime}_{k}}$. It verifies $\pi_{C}$; if invalid, aborts.
  10. (10)
    It extracts $\theta_{C}$ and $\overrightarrow{y^{\prime}}$ openings from the ZKP, sends $\theta_{C}$ to the trusted party, and receives score $\phi$.
  11. (11)
    Finally, it checks $\mathsf{Open}(\mathsf{pp},\mathtt{com}_{y^{\prime}_{i}},y^{\prime}_{i},r_{y^{\prime}_{i}})$ for each $i$. If any fail, it invokes $\mathcal{S}^{\mathsf{Score}}_{\mathsf{Alice}}$ for $\Pi_{\mathsf{SubScore}}$ with $\mathsf{abort}$, otherwise with $\phi$. It outputs whatever $\mathcal{S}^{\mathsf{Score}}_{\mathsf{Alice}}$ outputs.
The hybrids are:
  • $\mathcal{H}^{a}_{0}$: Real-world execution.
  • $\mathcal{H}^{a}_{1}$: Messages for $\Pi_{\mathsf{SubScore}}$ are simulated by $\mathcal{S}^{\mathsf{Score}}_{\mathsf{Alice}}$. Transition by security of $\Pi_{\mathsf{SubScore}}$.
  • $\mathcal{H}^{a}_{2}$: ZKP for model B inference replaced by $\mathcal{S}^{\prime}_{\mathsf{ZK}}$. Transition by zero-knowledge.
  • $\mathcal{H}^{a}_{3}$: Activations from model $a$ replaced by 0. Transition by hiding.
  • $\mathcal{H}^{a}_{4}$: Messages for $\Pi_{\mathsf{Inference}}$ simulated by $\mathcal{S}^{\mathsf{Inf}}_{\mathsf{Alice}}$. Transition by security of $\Pi_{\mathsf{Inference}}$.
  • $\mathcal{H}^{a}_{5}$: ZKP for challenge protocol $\mathsf{CP}$ simulated by $\mathcal{S}_{\mathsf{ZK}}$. Transition by zero-knowledge.
  • $\mathcal{H}^{a}_{6}$: Identical to ideal world except data points $(x,y)$ replaced by 0. Transition by hiding.
Transitions are justified as above. Thus the adversary’s view is indistinguishable from the ideal execution. This completes the proof that PrivaDE securely computes $\mathcal{F}_{\mathsf{Score}}$.
Appendix D Model Definitions
Here we show the exact model definitions of the distilled model in our experiment and the A-B-C splits. The split between B–C is evaluated with a data reconstruction attack, requiring the average SSIM score to be below 0.3.
Table 7 shows the LeNetXS model (used with MNIST dataset), table 8 shows the LeNet5 model (used with CIFAR-10) and table 9 shows the 5-layer CNN model (used with CIFAR-100).
Table 7. LeNetXS architecture with split bands.
# Layer (type) Output Shape Param #
[…]
Table 8. LeNet5 architecture with split bands.
# Layer (type) Output Shape Param #
[…]
Table 9. 5-layer CNN architecture with split bands.
# Layer (type) Output Shape Param #
[…]